A Complete Bibliography of ACM Transactions on Cyber-Physical Systems (TCPS)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

20 August 2024 Version 1.05

Title word cross-reference

1 [MGH^{+}23]. β [RRK^{+}22]. K [LHS^{+}21].

-Means [LHS^{+}21]. -Simplex [MGH^{+}23]. -VAE [RRK^{+}22].

1 [CJSM23, HZJ21].

2 [CJSM24, HZJ22]. 2021 [AO22].

4.0 [BMS^{+}22].

ABE [LYWY19]. Abnormal [WSH^{+}24]. Access [GVM^{+}20, WWW20]. Accounting [PM19]. Accuracy [WGH^{+}17]. across [GS18, MLW20]. Action [NRB^{+}18].

Action-Based [NRB^{+}18]. Actions [BGF22]. Activities [PRM^{+}20]. AcTrak [FPK^{+}23]. Adaptation [GBH^{+}17, SBW20]. Adaptive [CKP^{+}22, HFAD^{+}24, PPK18, TKP21, XBW^{+}19, ZMK^{+}17]. Additive [MVS^{+}19]. Advanced [WZZ^{+}20]. Adversarial [CCTS20, JHZ^{+}21]. Adversary [BRBB23]. Adversary-in-the-Loop [BRBB23]. Aerial [LLZ^{+}21]. Affine [YBKJ22]. Against [CLW^{+}24, ASGB23, CKT^{+}23, LTTY18, MLG24, ZDY^{+}20]. Agent [VLPB23]. Aggregation [APB^{+}21, YWZ^{+}19]. Agnostic [CLW^{+}24]. Agreement [LNE^{+}20]. Air [THR20]. AirTight [HBD^{+}20]. Alerts [BGGD20]. Algebra [BGF22]. Algorithm [WKS18]. Algorithmic [SBA21]. Algorithms [KRG17]. Allocation [PMKD22]. Analyses [SNBR23]. Analysing [BH21]. Analysis [CZV^{+}19, JZL^{+}17, KMM^{+}21, KAR22, KDM^{+}23, LLC^{+}19, LRN^{+}19, PKF24, THR20, WKS18, ZZC^{+}24, ZLSZ20]. Analytic [WZS^{+}19]. Analytics [GS18, WZR^{+}19]. AND/OR [BH21]. Android [WTZ^{+}19]. Anomaly [FBS^{+}21, LHS^{+}21, LCK^{+}19, PSM24, WZZ^{+}20]. Anonymous [TBKJ19]. Anti [KS22]. Anti-Lock [KS22]. Appliance [CKP^{+}22]. Application [BZS22, JHZ^{+}21]. Application-Driven [JHZ^{+}21]. Applications [ByKLS19, BGGT21, SST^{+}23, ST18, WWL^{+}19]. Approach [AR22, AK20, BMS^{+}22, CZV^{+}19, CKP^{+}22, KLMS20, LLC^{+}19, MJA^{+}23, MM22]. Apps [WTZ^{+}19]. Architecture [BDL^{+}24, DHS^{+}24, DFK^{+}24, DN23, KFH^{+}24, WHS18]. Architectures [BZS22]. Area [BTF^{+}21, PSM24]. Arrangements [MM22]. Arrays [FBS^{+}21]. Artificial [HZJ21, HZJ22]. ASHRAY [AEA^{+}22]. Aspects [AY17, AAH^{+}19]. Assessing [KOTD21]. Assessment [KAR22]. Assessments [AWB^{+}23]. Asset [MJA^{+}23]. Asset-driven [MJA^{+}23]. Assignment [GP20]. Assistance [WZZ^{+}20]. Assumptions [FGZ^{+}20]. Assurance [MJA^{+}23, TNY17]. Attack [ADD^{+}20, CLW^{+}24, KFH^{+}24, SST^{+}23, YZF18, ZZC^{+}24, WEL^{+}22]. Attacker [BGF22]. Attacks [GKMA22, KS22, KOTD21, LTTY18, LLC^{+}19, MLG24, PMAM21, YZF18, ZDY^{+}20, ASGB23]. Attestation [RKAG21]. Attitude [KMWB19]. Attribution [ADD^{+}20, WLTW22]. Augmenting [GKEQ21]. Authentication [ETM^{+}23, GVM^{+}20, TMF24, WWW20, YHL^{+}22]. Automata [YBKJ22]. Automated [AGG^{+}21, AR19, BRBB23, RZC^{+}22, SPH^{+}22]. Automatic [LCK^{+}19, WTZ^{+}19]. Automating [DN23]. Automation [BXL^{+}18, KMM^{+}21]. Automotive [CJSM23, CJSM24, CGCH18, CLW^{+}24, GKEQ21, KDM^{+}23, MJA^{+}23, PKF24, WLA19, XBW^{+}19, XZA^{+}19]. Autonomous [CBB22, KLMS20, KMA^{+}20, KMS20, LLZ^{+}21, MMTA24, SBA21, VLPB23, WWC22, YL23, YMH^{+}22]. Autonomy [VLPB23, WSH^{+}24]. Avionics [WEL^{+}22]. Avoidance [RPK^{+}21, YS20]. Aware [CGCH18, JLD20, LCC^{+}19, MMTA24, QXZ^{+}24, RPK^{+}21, XZA^{+}19, YWM^{+}19, BGGD20, BLD^{+}20, CZSL20, EB20, HMBP22, KLMS20, RP21, WZR^{+}19, XBW^{+}19]. awareness [MLW20].

B [KMA^{+}22]. Balancing [MHP^{+}21].

Bandwidth [MBH^{+}22]. Based

[AAD^{+}21, BCTV19, DAM^{+}18, HTU22, HSL^{+}19, JZL^{+}17, LYWY19, LZZD24, LCK^{+}19, LRN^{+}19, MN21, ANC^{+}23, MYK^{+}22, NRB^{+}18, PSW^{+}21, PSM24, RPK^{+}21, STB^{+}18, WKS18, WZZ^{+}20, XCL^{+}19, AGG^{+}21, BGGD20, GP20, ITB^{+}24, KMM^{+}21, KSK20, LNE^{+}20, LYA20, MLG24, MVS^{+}19, RP21, SZCL23, SNBR23, WZS^{+}19, WJL^{+}20, YFS^{+}19, YKDL24]. Battery [HKS17, HMWZ18]. Battery-Free [HMWZ18]. Bayesian [GWO^{+}20, HTU22]. Behavior [LCC^{+}19, XCL^{+}19]. Behavioral [BAU20]. Behaviors [WSH^{+}24, ZZC^{+}24]. between [HMWZ18]. Beyond [HMBP22, MBH^{+}22]. Bicycle [CLL^{+}21]. Bicycle-sharing [CLL^{+}21]. Big [LHS^{+}21, ZYCL19]. Black [TLW^{+}19]. Blockchain [AAD^{+}21, KFH^{+}24]. BloothAir [LLZ^{+}21]. BLSTM [LZZD24]. BLSTM-Based [LZZD24]. Bluetooth [LLZ^{+}21]. Bound [MYL^{+}20]. bounded [WAH22]. Box [TLW^{+}19]. Braking [KS22, MM22]. Build [MJA^{+}23]. Building [BKIS19, NVS^{+}21]. BuildingRules [NRB^{+}18]. Buildings [AR22, NRB^{+}18]. Bus [KS22, TdAGI^{+}22].

CACC [FCQY23]. CAD [KDM^{+}23]. Calibration [ZHZ19]. Camera [FPK^{+}23, MLG24]. Camera-based [MLG24]. Can [PMAM21, KS22, MRL^{+}23, TMF24]. CANOA [TMF24]. CANOPY [PSW^{+}21]. Capacity [HKS17]. Cardiac [YAR^{+}18]. Carriers [FGZ^{+}20]. CASCADE [MJA^{+}23]. Case [BAU20, GGB^{+}18, HKS17, TSC21]. Cases [MJA^{+}23]. CASTNet [MMTA24]. Categorical [BFV21]. Catering [WTZ^{+}19]. CAV [FWL23]. Center [WLTW22, WCZ^{+}24]. centeric [APB^{+}21]. Centralized [KMM^{+}21]. Centric [SCB^{+}20, WTZ^{+}19, AK20, HZW^{+}20, PYD^{+}20]. Challenges [JHZ^{+}21, MRB^{+}22, VBV^{+}19]. Changes [BMJ^{+}20]. channel [BZW^{+}20, TMF24]. Channels [MH23]. Characteristics [ARC^{+}23]. Characterization [WD21]. Charging [BCDT24, GKMA22, SST^{+}23, YS21]. CHARIOT [PDK^{+}18]. CirclePIN [GVM^{+}20]. Cities [MPA^{+}20, TYT^{+}23]. City [YMH^{+}22]. Civilian [AY17]. Class [GKMA22]. Classification [CCTS20]. Classify [PLLH22]. Clock [RTY^{+}19]. Close [MM22, YWM^{+}19]. Close-Distance [MM22]. Closed [PPK18]. Closed-Loop [PPK18]. Cloud [GS18, YNL^{+}22]. Clustering [ADD^{+}20]. Co [BTF^{+}21, DN23, FCQY23, NVS^{+}21, WLTW22]. Co-Design [BTF^{+}21, DN23, FCQY23]. Co-Located [WLTW22]. Co-working [NVS^{+}21]. Code [KFSL19]. Cognitive [KAR22]. Collaborative [BTB^{+}18, FWL23, HTU22, TNY17]. Collection [AAD^{+}21]. Collectives [EB20]. Collision [RPK^{+}21]. Combinatorial [ZZC^{+}24]. Combined [GWO^{+}20, MH23]. Combined-Slip [MH23]. Combining [Asp20]. Comfort [AEA^{+}22]. Commercial [NRB^{+}18]. Commodity [XCL^{+}19]. Commonalities [SNBR23].

Communication

[BZW^{+}20, BTF^{+}21, FCQY23, LYWY19, MYL^{+}20, PSW^{+}21, WLA19, ZMK^{+}17]. Communications [GRA^{+}21, MH23]. Communities [PMKD22]. Comparison [PSM24, QXZ^{+}24]. Compilation [BGGT21]. Complex [NRB^{+}18]. Component [LSN^{+}18]. Composable [RP21]. Compression [HHS19]. Compromise [ASGB23]. Compromised [BAU20]. Computation [HSL^{+}19, XZA^{+}19, ZYCL19]. Computing [SZCL23, SJT^{+}22, YWZ^{+}19]. Concepts [DHS^{+}24]. Concerns [WTZ^{+}19]. Concurrent [TdAGI^{+}22]. Conduction [YAR^{+}18]. Confidence [BXL^{+}18]. Configuration [LYA20]. Configurations [MN21]. Conflict [YMH^{+}22]. Conformal [QXZ^{+}24]. Conformance [ROWA19]. Congestion [YS20]. Connected [ACLY21, KLMS20, KMA^{+}20, KMS20, LLZ^{+}21, PMKD22, RSJ21, WWC22]. Conquer [PLLH22]. Considered [WWC22]. Consistently [MMSM21]. Constant [FWL23]. Constant-Time-Headway [FWL23]. Constrained [LJP20, PM19, TSC21]. Constraint [ANC^{+}23]. Constraints [HCT^{+}23]. Containers [CKT^{+}23]. Contaminated [LZZD24]. Context [MMTA24]. Context-Aware [MMTA24]. Contiki [MEM23]. Continuous [YHL^{+}22]. Control [ASGB23, BBCV20, BBCV21, BMJ^{+}20, BTF^{+}21, BZS22, CBB22, ETM^{+}23, HMBP22, HFAD^{+}24, KFH^{+}24, KMWB19, MGL^{+}19, MLW20, MBH^{+}22, MH23, MMSM21, PKF24, SLJ^{+}23, TLW^{+}19, TdAGI^{+}22, VTK24, WHS18, WCZ^{+}24]. Controller [CGCH18, MVS^{+}19, PSM24]. Controllers [MLW20, RZC^{+}22]. Controlling [FPK^{+}23]. Convergence [PYD^{+}20]. Convolutional [TSC21]. Cooling [NVS^{+}21, WLTW22, WCZ^{+}24]. Cooperating [WWC22]. Cooperation [Asp20]. Cooperative [HFAD^{+}24]. Coordinated [GKMA22]. Coordination [VMDJ20]. copter [HTU22]. Core [CKT^{+}23]. Corner [YL23]. Cost [XZA^{+}19]. COTS [ZMC^{+}23]. Countermeasures [LTT^{+}20, MLG24]. CPS [SD21, BAU20, BGGD20, BDL^{+}24, CJSM23, CJSM24, GGB^{+}18, HBD^{+}20, JSY^{+}19, SCB^{+}20, TLW^{+}19, WZZ^{+}20]. Critical [LSN^{+}18, LJZ18, SNBR23, BH21, KDM^{+}23]. Criticality [HBD^{+}20, GP20]. Cross [TLW^{+}19]. Cross-Domain [TLW^{+}19]. Crossing [MM23]. Crossroads [KLMS20]. Crowd [APB^{+}21]. Crowd-sensed [APB^{+}21]. Crowdsensing [LSN^{+}18, YFS^{+}19]. Crowdsensing-based [YFS^{+}19]. Cruise [HFAD^{+}24]. CSIP [AR19]. Cyber [ADD^{+}20, AAH^{+}19, ACLY21, ASGB23, BFV21, BBCV20, BRBB23, BH21, BLD^{+}20, BDL^{+}24, BTF^{+}21, ByKLS19, CZSL20, CH20, DHS^{+}24, DFK^{+}24, EB20, FGZ^{+}20, GP20, GBH^{+}17, HZJ21, HZJ22, HHS19, JAR^{+}20, JLD20, JRGB^{+}18, KOTD21, KMA^{+}22, KAR22, KSK20, Lee17, LP18, LJP20, LCC^{+}19, LNE^{+}20, LCK^{+}19, LYA20, Lu21, MGL^{+}19, ANC^{+}23, MRB^{+}22, MYL^{+}20, MM22, NHB^{+}18, PLLH22, PKF24, PIW^{+}17, PRM^{+}20, RRK^{+}22, ROWA19, RP21, SJT^{+}22, SSV^{+}18, SS22, SDZV19, Sta17, THR20, WLA19, WGH^{+}17, WMC18, WHS18, WZS^{+}19, WYW^{+}19, WWL^{+}19, WGL19, WZR^{+}19, WEL^{+}22, XCL^{+}19, XBW^{+}19, XZA^{+}19, YFS^{+}19, YKDL24, YZF18, YWZ^{+}19, ZYCL19, ZA18a, ZA18b, ZJP21, ZMK^{+}17, ZGAB21]. Cyber-attack [WEL^{+}22]. Cyber-attacks [ASGB23]. Cyber-Physical [ADD^{+}20, AAH^{+}19, BFV21, BBCV20, BRBB23, BLD^{+}20, BTF^{+}21, ByKLS19, CH20, DHS^{+}24, DFK^{+}24, FGZ^{+}20, GBH^{+}17, HZJ21, HZJ22, HHS19, JAR^{+}20, JRGB^{+}18, KOTD21, KMA^{+}22, Lee17, LP18, LJP20, LCC^{+}19, LCK^{+}19, Lu21, MGL^{+}19, ANC^{+}23, MRB^{+}22, MYL^{+}20, NHB^{+}18, PIW^{+}17, RRK^{+}22, ROWA19, SSV^{+}18, SDZV19, WLA19, WMC18, WZS^{+}19, WWL^{+}19, WGL19, XCL^{+}19, XBW^{+}19, XZA^{+}19, YKDL24, YZF18, YWZ^{+}19, ZYCL19, ZA18a, ZA18b, ZMK^{+}17, ZGAB21, ACLY21, BH21, CZSL20, EB20, GP20, JLD20, KAR22, KSK20, LNE^{+}20, LYA20, MM22, PKF24, PRM^{+}20, RP21, SJT^{+}22, SS22, WGH^{+}17, WZR^{+}19, YFS^{+}19, ZJP21]. Cyber-Physical-Social [WYW^{+}19]. Cyber-risk [PLLH22]. Cybersecurity [MRB^{+}22].

Daily [PRM^{+}20]. Dashboard [ZMC^{+}23]. Data [AEA^{+}22, AAD^{+}21, GWO^{+}20, HHS19, HZW^{+}20, KMM^{+}21, KRG17, LTTY18, LHS^{+}21, LYWY19, LZZD24, LTT^{+}20, MPA^{+}20, MHP^{+}21, WMC18, WWW20, WLTW22, WCZ^{+}24, WZR^{+}19, YZF18, YWZ^{+}19, YL23, ZZZ^{+}17a, ZZZ^{+}17b, ZYCL19, ZHZ19, ZDY^{+}20]. Data-Driven [YL23, AEA^{+}22, MHP^{+}21]. Decentralized [RKAG21, RPK^{+}21, VLPB23, YMH^{+}22]. Deception [BBCV21]. Deception-Robust [BBCV21]. Decision [BGGD20, BB21, MPA^{+}20, MRL^{+}23]. Decomposition [HHS19, WYW^{+}19]. Deep [AGG^{+}21, HFAD^{+}24, MN21, SS22, TKP21, ZYCL19]. Deep-Learning [TKP21]. Defense [BRBB23, KSK20, ZDY^{+}20]. Delay [HMBP22, HFAD^{+}24, ZLSZ20]. Delay-aware [HMBP22]. Delay-Tolerant [ZLSZ20]. Delivery [HKS17]. Demand [MHP^{+}21]. Dense [LRN^{+}19]. Dependability [ByKLS19, WWL^{+}19]. Dependable [JCO21, JSY^{+}19, LCK^{+}19, MGL^{+}19, WZS^{+}19, ZYCL19]. Deploying [AR22]. Deployments [BBS21]. DeResolver [YMH^{+}22]. Derivation [LYA20]. Descriptions [WTZ^{+}19]. Design [BDL^{+}24, BTF^{+}21, CGCH18, FGZ^{+}20, HSL^{+}19, KMA^{+}22, LNE^{+}20, NC21, WZ17, XZA^{+}19, ZLSZ20, DN23, FCQY23]. Designed [PMAM21]. Designing [ANC^{+}23, MVS^{+}19]. Detecting [LTTY18, WSH^{+}24, ZZC^{+}24]. Detection [ARC^{+}23, Asp20, BAU20, CZSL20, DYKE24, DAM^{+}18, FBS^{+}21, GRA^{+}21, ITB^{+}24, JCO21, KS22, LHS^{+}21, LZZD24, LCK^{+}19, MSZ^{+}23, MRL^{+}23, RRK^{+}22, SZCL23, WJL^{+}20, WEL^{+}22, WZZ^{+}20, YKDL24, ZGAB21]. Detectors [PSM24]. Determining [KFSL19]. Developing [AEA^{+}22, TSC21]. Development [JSY^{+}19]. Device [STB^{+}18]. Devices [BAU20, GVM^{+}20, HMWZ18, MYL^{+}21, PXH^{+}19, TKP21, TBKJ19, TSC21, ZMC^{+}23]. Diagnosis [GRA^{+}21]. Differences [SNBR23]. Differentially [NC21]. Dimension [XCL^{+}19]. Directions [Sta17]. Discharging [GKMA22]. Discipline [Lu21]. Dispatching [CLL^{+}21]. Disruptions [BMS^{+}22]. Distance [MM22]. Distinct [WLTW22]. Distributed [GKEQ21, GS18, IIL^{+}17, MSZ^{+}23, RSJ21, RZC^{+}22, STB^{+}18, TBKJ19, WYW^{+}19]. Distribution [BCTV19, DYKE24, PMAM21, RRK^{+}22, YKDL24]. Distributionally [MHP^{+}21]. Dockless [CLL^{+}21]. Domain [TLW^{+}19, KAR22]. Driven [BTB^{+}18, JHZ^{+}21, KRG17, PDK^{+}18, YL23, AEA^{+}22, FBS^{+}21, JSY^{+}19, MHP^{+}21, MJA^{+}23]. Driver [WZZ^{+}20]. Drivers [WSH^{+}24]. Driving [AGG^{+}21, BZW^{+}20, BHX^{+}20, MGH^{+}23, MMTA24, MM22, YS20, YL23]. Drone [YFS^{+}19]. Drones [ARC^{+}23, AY17, LLZ^{+}21]. Drop [BGGD20]. DS [MRL^{+}23]. DT [MRL^{+}23]. DT-DS [MRL^{+}23]. Dumb [IKG^{+}19]. Duty [MVS^{+}20, VMDJ20]. Dynamic [CLL^{+}21, GAT20, KSK20, MGH^{+}23, MMTA24, PMKD22, THR20, VMDJ20, WKS18].

ECG [HHS19]. Edge [ADD^{+}20, GS18, PXH^{+}19, PYD^{+}20, STB^{+}18, SJT^{+}22].

Editorial [LP18]. Effect [MH23]. Effective [ZA18a, ZA18b]. Effects [KFH^{+}24]. Efficiency [YS21]. Efficient [HHS19, MLW20, MYL^{+}21, MYK^{+}22, PKF24, RRK^{+}22, SSV^{+}18, TdAGI^{+}22, WWW20, YNL^{+}22, YL23, ZA18a, ZA18b]. Electric [BCDT24, GKMA22, KRG17, YS21]. Electricity [GWO^{+}20, LZZD24]. Embedded [BGGT21, RZC^{+}22]. Emergence [CZV^{+}19]. Emergence-Oriented [CZV^{+}19]. Emergency [MM22]. Emphasizing [SJT^{+}22]. Empirical [HHS19]. Emulation [BBS21, YAR^{+}18]. Enabled [YKDL24]. Enabling [IIL^{+}17, TdAGI^{+}22]. Encrypted [YNL^{+}22]. End [XZA^{+}19]. End-to-End [XZA^{+}19]. Energy [HZI^{+}17, HHS19, IIL^{+}17, LCC^{+}19, MYK^{+}22, NVS^{+}21, SSV^{+}18, SLJ^{+}23]. Energy-Efficient [HHS19, MYK^{+}22, SSV^{+}18]. Engineering [Lu21]. Enhance [WGH^{+}17]. Enhancing [AEA^{+}22, BZW^{+}20, VTK24]. Ensemble [MMTA24]. Ensembles [MRL^{+}23]. Ensuring [BXL^{+}18]. Entry [CLW^{+}24]. Environment [BBS21, GRA^{+}21]. Environmental [JCO21, TSC21]. Environments [HTU22, KMS20, MGH^{+}23]. Erasure [MH23]. Estimation [YZF18]. Ethernet [MN21]. Ethics [SBA21]. EV [SST^{+}23]. Evaluation [BCTV19, CBB22, DAM^{+}18, TLW^{+}19]. Event [BGGD20, CZSL20, FCQY23, GS18, JZL^{+}17, MH23, TdAGI^{+}22, WMC18, WGL19]. Event-Based [JZL^{+}17]. Event-Triggered [FCQY23, MH23, TdAGI^{+}22]. EVScout2.0 [BCDT24]. Examples [JHZ^{+}21]. Execution [GKEQ21]. Experimental [SSV^{+}18]. experts [CKP^{+}22]. Exploiting [WGH^{+}17]. Exploration [HTU22]. Explosion [BVS^{+}21]. Extending [LHS^{+}21, ZJP21]. Extensive [LRN^{+}19].

Factor [WWW20]. Failures [KMA^{+}22, WHS18]. False [LTTY18, ZDY^{+}20]. Fast [BMJ^{+}20, TKP21]. Fault [JAR^{+}20, TBKJ19, XZA^{+}19]. Fault-Tolerant [TBKJ19, XZA^{+}19]. Faults [PIW^{+}17]. Feasibility [MN21]. Feature [BTB^{+}18, ZYCL19]. Feature-Driven [BTB^{+}18]. Federated [BMS^{+}22]. Feedback [BMJ^{+}20]. Feet [HMWZ18]. Filter [HFAD^{+}24]. Finite [RP21, YFS^{+}19]. Firmware [RKAG21]. Flexible [BVS^{+}21]. Flight [MMSM21]. Flow [VLPB23]. Flows [THR20]. Fluctuation [WD21]. Flying [YWM^{+}19, RPK^{+}21]. Fog [BMS^{+}22]. Fog-supported [BMS^{+}22]. Follower [MH23]. Forecasting [GWO^{+}20]. Forest [LHS^{+}21]. Formal [BZS22, GGB^{+}18, MEM23, PKF24, WJL^{+}20]. Formalization [SBA21]. Formulae [BB21]. Forward [EEC^{+}21, WWW20]. Forward-trading [EEC^{+}21]. Foundations [DFK^{+}24]. FPV [ARC^{+}23]. Framework [ARC^{+}23, BAU20, GAT20, HSL^{+}19, ANC^{+}23, MMSM21, SS22, TSC21, VMDJ20, WZS^{+}19, YBKJ22, YMH^{+}22]. Free [HMWZ18]. FreeSia [KAR22]. Frequency [AK20, KAR22]. Frequency-domain [KAR22]. Frequency-regulating [AK20]. Full [JAR^{+}20]. Functional [XBW^{+}19, XZA^{+}19]. Fundamental [BDL^{+}24, DHS^{+}24, Lee17]. Fuzzy [ADD^{+}20].

Game [MYK^{+}22, YS21]. Gas [BCTV19, MRB^{+}22]. Gateway [STB^{+}18]. Gaussian [GWO^{+}20]. Generation [KFSL19, KMS20, MYL^{+}20, WLA19, WTZ^{+}19]. Generators [AK20]. Goal [PDK^{+}18]. Goal-Driven [PDK^{+}18]. Gordian [WJL^{+}20]. Graphs [BH21]. Green [WCZ^{+}24]. Grid [BAU20, RTY^{+}19, SZCL23, WZS^{+}19, ZDY^{+}20]. Grids [KRG17, LLC^{+}19, TNY17]. Groups [Asp20]. Guarantee [FWL23]. Guarantees [BMJ^{+}20, MBH^{+}22, YKDL24]. Guest [LP18]. guided [WCZ^{+}24].

Hacking [BGF22]. Hand [BHX^{+}20]. Hard [HCT^{+}23, PLLH22]. Hardness [PLLH22]. Harvesting [PXH^{+}19, SLJ^{+}23]. Headway [FWL23]. Health [DAM^{+}18, MPP^{+}24]. Healthcare [Sta17]. Heavy [MVS^{+}20, VMDJ20]. Heavy-Duty [VMDJ20]. Heterogeneous [ZZZ^{+}17a]. Hidden [YHL^{+}22]. Hiddenness [ZDY^{+}20]. Hierarchical [PMKD22, TYT^{+}23]. Highway [FWL23]. Holistic [MGL^{+}19, MLW20]. Home [BXL^{+}18, KMM^{+}21, LCC^{+}19, PRM^{+}20]. Homes [CKP^{+}22, PSW^{+}21]. Hood [BGMM19]. Hop [TdAGI^{+}22, BMJ^{+}20]. HPRoP [TYT^{+}23]. HUCDO [HZW^{+}20]. Hue [KMM^{+}21]. Human [BDL^{+}24, DHS^{+}24, DFK^{+}24, SPH^{+}22, WGH^{+}17, WSH^{+}24, WZR^{+}19, XBW^{+}19, BDL^{+}24]. Human-CPS [BDL^{+}24]. Human-Interaction-aware [XBW^{+}19, WZR^{+}19]. Humans [YWM^{+}19]. Humidities [WLTW22]. Hybrid [HZW^{+}20, SS22, WAH22, WSH^{+}24, YBKJ22]. Hypergames [BBCV20, BBCV21].

ICCPS [AO22, MV24]. ICPS [MRB^{+}22]. ICSs [PLLH22]. Identification [YBKJ22]. II [BDL^{+}24, ZA18b]. III [DFK^{+}24]. Image [MVS^{+}19]. Image-based [MVS^{+}19]. Impact [THR20, TLW^{+}19]. Impairment [DAM^{+}18]. Implementation [JSY^{+}19, LNE^{+}20]. Imprecise [HSL^{+}19]. Improve [JRGB^{+}18, WWL^{+}19, ZGAB21]. Improved [XCL^{+}19]. Improving [BHX^{+}20, HKS17, VBV^{+}19]. In-motion [YS21]. In-Vehicle [RKAG21]. Incentive [NC21]. Incentivizing [IIL^{+}17]. Incident [ITB^{+}24]. Incision [GGB^{+}18]. Incomplete [ZHZ19]. Increase [KFH^{+}24]. Independent [KFSL19]. Indicators [ASGB23]. Indoor [KAR22, LCK^{+}19, TKP21]. Industrial [ASGB23, ETM^{+}23, KOTD21, KFH^{+}24, MRB^{+}22, VTK24, ZZC^{+}24]. Industry [MRB^{+}22, BMS^{+}22]. Inference [QXZ^{+}24]. Inferring [IKG^{+}19]. Information [RP21, WSH^{+}24]. Infrastructure [ZZZ^{+}17a, ZZZ^{+}17b]. Initial [KFH^{+}24]. Injection [LTTY18, YZF18, ZDY^{+}20]. Input [HHS19, YBKJ22]. Input-Output [YBKJ22]. Insuring [PLLH22]. Integrating [LJP20]. Integration [HSI^{+}20, RZC^{+}22, ZZZ^{+}17a]. Integrity [LTT^{+}20]. Intelligence [HZJ21, HZJ22]. Intelligent [MM23, SS22, ZLSZ20]. Interaction [BDL^{+}24, EB20, XBW^{+}19, WZR^{+}19]. Interactions [WGH^{+}17]. Interactive [CKP^{+}22, LYA20]. Intermittent [MH23]. Internet [APB^{+}21, BTB^{+}18, DRC^{+}18, LSN^{+}18, MSZ^{+}23, PYD^{+}20, WZ17, ZA18a, ZA18b, ZZC^{+}24]. Interpretable [DYKE24]. Intersection [KLMS20, KMA^{+}20, WWC22, ZLSZ20]. Intersections [AR19, MM23]. Intervals [ZJP21]. Introduction [AO22, AAH^{+}19, ACLY21, CH20, CJSM23, CJSM24, HZJ21, HZJ22, JLD20, Kuo17, MV24, SCB^{+}20, SD21, WZR^{+}19]. Introspection [BGMM19]. Intrusion [ARC^{+}23, GRA^{+}21, MRL^{+}23, SZCL23, WEL^{+}22, ZGAB21]. Invasive [ARC^{+}23]. Investigating [SST^{+}23]. IoT [AEA^{+}22, AAD^{+}21, BBS21, BXL^{+}18, GVM^{+}20, PXH^{+}19, PDK^{+}18, STB^{+}18, ST18, WZS^{+}19]. IoT-based [WZS^{+}19]. IoTranx [CHJ^{+}22]. Isolation [LHS^{+}21]. Issue [AAH^{+}19, ACLY21, CH20, CJSM23, CJSM24, HZJ21, HZJ22, JLD20, LP18, MV24, SCB^{+}20, SD21, WZR^{+}19, ZA18a, ZA18b]. IT/OT [PLLH22].

Just [HMBP22].

Key [LNE^{+}20, WLA19]. Keyless [CLW^{+}24]. Knowledge [HSI^{+}20]. KP [LYWY19]. KP-ABE [LYWY19].

Large [BBS21, JRGB^{+}18, SDZV19].

Large-Scale [SDZV19]. Laser [GGB^{+}18]. Last [ZZZ^{+}17b]. Last-Mile [ZZZ^{+}17b]. Latency [MYL^{+}20, BMS^{+}22]. Latent [DYKE24, RRK^{+}22]. Layer [ADD^{+}20, WLA19]. LDA [XCL^{+}19]. Leader [MH23]. Leader-Follower [MH23]. Leaks [CKT^{+}23]. Learning [AGG^{+}21, BB21, BMS^{+}22, CLL^{+}21, CKP^{+}22, FPK^{+}23, HSL^{+}19, JCO21, KMWB19, MN21, MMSM21, RPK^{+}21, SZCL23, SBW20, SS22, TKP21, VLPB23, WCZ^{+}24, XCL^{+}19, YKDL24, ZYCL19, ZGAB21, RPK^{+}21]. Learning-Based [HSL^{+}19, RPK^{+}21, AGG^{+}21, SZCL23]. Learning-‘N-Flying [RPK^{+}21]. Ledger [RSJ21]. Level [AWB^{+}23, TLW^{+}19, BAU20, GKEQ21]. Leveraging [HHS19]. Light [LCK^{+}19]. Light-Based [LCK^{+}19]. Limb [BGM^{+}18]. Limitations [MBH^{+}22]. Limits [Lee17]. Lines [LYA20]. Link [WD21]. Linux [CKT^{+}23]. Load [BKIS19, GWO^{+}20, IIL^{+}17, TNY17]. Localization [LCK^{+}19, MSS18, TKP21, WWL^{+}19, WJL^{+}20]. Located [WLTW22]. Location [LCK^{+}19]. Location-Based [LCK^{+}19]. Lock [KS22]. Locomotion [KAR22]. Logic [BB21, HTU22, ZJP21]. Logical [GKEQ21]. Long [WMC18]. Long-Term [WMC18]. Looking [BGMM19]. Loop [BRBB23, PPK18]. Lord [CBB22]. Loss [HZI^{+}17]. Low [BMS^{+}22, PXH^{+}19]. Low-latency [BMS^{+}22]. LSTM [WZZ^{+}20]. Lustre [BGGT21].

Machine [JCO21, RP21, SZCL23]. Machine-based [RP21]. Maintaining [CZSL20]. Maintenance [SS22, ZMC^{+}23].

Makespan [BVS^{+}21]. Making [BGGD20, MPA^{+}20]. Malicious [ZZC^{+}24]. Malware [ADD^{+}20]. Manage [NRB^{+}18]. Management [CZSL20, JRGB^{+}18, KLMS20, KMA^{+}20, LCC^{+}19, MGL^{+}19, MSZ^{+}23, PLLH22, STB^{+}18, TNY17, WWC22, ZLSZ20]. Manufacturing [BVS^{+}21]. Mapping [SJT^{+}22]. Markov [YHL^{+}22]. MaxSAT [BH21]. MC [BZW^{+}20]. MC-Safe [BZW^{+}20]. Mean [ITB^{+}24]. Mean-based [ITB^{+}24]. Means [LHS^{+}21]. Mechanism [GVM^{+}20]. Mediator [BTB^{+}18]. Medical [FGZ^{+}20, GGB^{+}18, LP18]. Medium [MVS^{+}20]. Memory [BGMM19, YKDL24]. Memory-based [YKDL24]. Merging [FWL23]. Message [ETM^{+}23, RSJ21]. Meta [GBH^{+}17, ZGAB21]. Meta-Adaptation [GBH^{+}17]. Meta-Learning [ZGAB21]. Metagames [BBCV21]. Meteorological [DYKE24]. Metered [FWL23]. Metered-Ramp [FWL23]. Meters [BKIS19]. Method [WYW^{+}19, ZZC^{+}24]. Methodology [AAD^{+}21, KMA^{+}22, SSV^{+}18, XZA^{+}19]. Metrics [CCTS20]. Microgrids [EEC^{+}21, HZI^{+}17]. Middleware [PDK^{+}18, ST18, WGL19]. Mile [ZZZ^{+}17b]. Minimal [SNBR23]. Minimization [YS20]. Minimizing [HZI^{+}17]. Minimum [AR22]. Mining [FGZ^{+}20]. Misbehavior [MSZ^{+}23]. Mismatches [NHB^{+}18]. Mission [BH21, RPK^{+}21]. Mission-Aware [RPK^{+}21]. Mission-critical [BH21]. Mitigation [KS22, YZF18]. Mixed [GP20, HBD^{+}20, VLPB23]. Mixed-criticality [GP20]. Mobile [DAM^{+}18, LSN^{+}18, LCK^{+}19, MYL^{+}21, SST^{+}23, Sta17, SEB21, TKP21, WD21, ZMC^{+}23]. MobileTrust [HSI^{+}20]. Mobility [GAT20, JZL^{+}17, MHP^{+}21]. Mobility-on-Demand [MHP^{+}21]. Mode [BMJ^{+}20, HHS19]. Model [BCTV19, FBS^{+}21, JSY^{+}19, ROWA19, SNBR23, WAH22, ZZZ^{+}17a, ZYCL19, ZHZ19]. Model-Based [BCTV19, SNBR23]. Model-bounded [WAH22]. Model-driven [FBS^{+}21, JSY^{+}19]. Modeling [AWB^{+}23, CCTS20, JZL^{+}17, LTTY18, Lee17, MPA^{+}20, PXH^{+}19, RP21, YZF18]. Modelling [LJZ18]. Models [EB20, FGZ^{+}20, KFSL19, QXZ^{+}24, YHL^{+}22]. Modern [MYK^{+}22]. Modularity [BZS22]. Monitoring [BMS^{+}22, HMBP22, HCT^{+}23, JCO21, MPP^{+}24, TMF24, TSC21, WAH22, ZJP21]. Motion [BHX^{+}20, MMTA24, YS21]. Movement [CCTS20]. Moving [ZDY^{+}20]. Multi [BZW^{+}20, BMJ^{+}20, CLL^{+}21, CKT^{+}23, STB^{+}18, TdAGI^{+}22, VLPB23, WWW20, XBW^{+}19, ZZZ^{+}17a, ZHZ19]. Multi-Agent [VLPB23]. Multi-channel [BZW^{+}20]. Multi-Core [CKT^{+}23]. Multi-Factor [WWW20]. Multi-Functional [XBW^{+}19]. Multi-Gateway [STB^{+}18]. Multi-Hop [TdAGI^{+}22, BMJ^{+}20]. Multi-objective [CLL^{+}21]. Multi-Source [ZZZ^{+}17a, ZHZ19]. Multilabel [ADD^{+}20]. Multiple [HCT^{+}23].

N [RPK^{+}21]. National [ZHZ19].

National-scale [ZHZ19]. Negotiation [YMH^{+}22]. Nest [KMM^{+}21]. Net [LLC^{+}19, WZ17]. Nets [LJZ18]. Network [LRN^{+}19, MN21, PSM24, WEL^{+}22, WZZ^{+}20]. Networking [LYWY19]. Networks [ADD^{+}20, APB^{+}21, BMJ^{+}20, BCTV19, MLW20, RKAG21, SEB21, THR20, TSC21, WD21]. Neural [TSC21, WZZ^{+}20]. Neurocognitive [DAM^{+}18]. Neuroevolutionary [HSL^{+}19]. Nodes [SBW20, SLJ^{+}23]. Noise [TLW^{+}19]. Non [CGCH18, CKP^{+}22, PXH^{+}19]. Non-experts [CKP^{+}22]. Non-Uniform [CGCH18]. Non-volatile [PXH^{+}19]. Novel [AR22, GVM^{+}20].

Object [MLG24]. objective [CLL^{+}21]. Obligations [SBA21]. OD1NF1ST [WEL^{+}22]. Offline [BB21]. Offload [NC21]. Offshore [MRB^{+}22]. Oil [MRB^{+}22]. Online [BKIS19, BB21]. Operating [JRGB^{+}18]. Operation [SSV^{+}18]. Operational [MVS^{+}20]. Operator [YHL^{+}22]. Optimal [NVS^{+}21]. Optimization [BVS^{+}21, MHP^{+}21, PXH^{+}19, ZZC^{+}24]. Optimize [YS21]. Optimizing [VLPB23, YS20]. Orchestration [PDK^{+}18]. Oriented [CZV^{+}19]. Origin [TMF24]. OS-Aware [CGCH18]. OT [PLLH22]. Out-of-Distribution [DYKE24, RRK^{+}22]. Outlier [JCO21, WJL^{+}20]. Output [YBKJ22]. Outsourcing [HZW^{+}20]. Overload [MBH^{+}22].

P300 [DAM^{+}18]. Pacemaker [YAR^{+}18]. Pacemakers [PPK18]. Packs [HKS17]. panel [FBS^{+}21]. Papers [AO22]. Paradigm [GKEQ21]. Parameter [MYK^{+}22]. Parameterized [YL23]. Parameters [KFSL19]. Part [BDL^{+}24, CJSM23, CJSM24, DHS^{+}24, DFK^{+}24, HZJ21, HZJ22, ZA18b]. Participatory [BGGD20]. Partnership [PRM^{+}20]. Path [HTU22, MYK^{+}22, YWM^{+}19]. Pattern [LYA20]. Pattern-based [LYA20]. Patterns [DN23]. Pedestrians [WWC22]. People [APB^{+}21]. Per-panel [FBS^{+}21]. Perception [AGG^{+}21, MLG24, YL23]. Performance [PKF24, PSM24]. Periodic [FCQY23]. Personalised [WTZ^{+}19]. Perspective [ASGB23, MRB^{+}22]. Pervasive [LYWY19]. Petri [LLC^{+}19, LJZ18]. Physical [ADD^{+}20, AAH^{+}19, BFV21, BBCV20, BRBB23, BLD^{+}20, BDL^{+}24, BTF^{+}21, ByKLS19, CH20, CCTS20, DHS^{+}24, DFK^{+}24, FGZ^{+}20, GBH^{+}17, HZJ21, HZJ22, HHS19, JAR^{+}20, JRGB^{+}18, KOTD21, KMA^{+}22, Lee17, LP18, LJP20, LCC^{+}19, LCK^{+}19, Lu21, MGL^{+}19, ANC^{+}23, MRB^{+}22, MYL^{+}20, NHB^{+}18, PIW^{+}17, RRK^{+}22, ROWA19, SSV^{+}18, SDZV19, Sta17, WLA19, WMC18, WHS18, WZS^{+}19, WYW^{+}19, WWL^{+}19, WGL19, XCL^{+}19, XBW^{+}19, XZA^{+}19, YKDL24, YZF18, YWZ^{+}19, ZYCL19, ZA18a, ZA18b, ZMK^{+}17, ZGAB21, ACLY21, BH21, CZSL20, EB20, GP20, JLD20, KAR22, KSK20, LNE^{+}20, LYA20, MM22, PKF24, PRM^{+}20, RP21, SJT^{+}22, SS22, WGH^{+}17, WZR^{+}19, YFS^{+}19, ZJP21]. Physics [WCZ^{+}24]. Physics-guided [WCZ^{+}24]. Pipelined [MVS^{+}19]. Placement [NVS^{+}21]. Plan [KMA^{+}22]. Planning [BRBB23, HTU22, MYK^{+}22, PMKD22, SPH^{+}22, TYT^{+}23, YWM^{+}19]. Platform [BGM^{+}18, EEC^{+}21, KFSL19]. Platform-Independent [KFSL19]. Platforms [KMM^{+}21, YNL^{+}22]. Platoon [LNE^{+}20, MM23, VMDJ20]. Platoon-based [LNE^{+}20]. Platooning [FCQY23]. Policies [CBB22]. Power [BKIS19, CZSL20, LTTY18, PXH^{+}19, RTY^{+}19, TMF24, WLTW22, WKS18, ZDY^{+}20]. powered [CZSL20, PXH^{+}19]. Predict [MN21]. Prediction [MMTA24]. Predictive [SS22, VMDJ20]. Preface [ZA18a, ZA18b]. Presence [PIW^{+}17]. Preserving [PSW^{+}21, SEB21, APB^{+}21, TYT^{+}23, YWZ^{+}19]. Prevent [GVM^{+}20]. Principles [BDL^{+}24]. Priority [MN21]. Priority-Based [MN21]. Privacy [AY17, ACLY21, APB^{+}21, DRC^{+}18, PSW^{+}21, TYT^{+}23, YWZ^{+}19]. Privacy-Preserving [PSW^{+}21, APB^{+}21, TYT^{+}23, YWZ^{+}19]. Private [EEC^{+}21, NC21]. Probability [CZSL20]. Problems [MH23]. Procedures [BCTV19]. Process [GWO^{+}20]. Processing [KRG17, WMC18, WGL19, XBW^{+}19]. Processors [CKT^{+}23]. Product [LYA20]. Profile [BCDT24]. Profiling [BCDT24].

Progress [JHZ^{+}21]. Propagation [AWB^{+}23]. Proposal [KFH^{+}24]. Protecting [CKT^{+}23]. Protocol [AR19, FWL23, HBD^{+}20, MM23, PSW^{+}21, WWW20]. Provenance [ETM^{+}23]. Proximity [YWM^{+}19]. Pruning [MYL^{+}21]. Public [CLL^{+}21, NC21]. Publish [BGGD20, ST18]. Publish/Subscribe [ST18]. Pythagorean [ITB^{+}24].

QoS [KRG17]. QoS-Driven [KRG17]. Quadrotor [MMSM21]. Quality [RP21, WD21, BGGD20]. Quality-aware [BGGD20]. Quality-of-Information-aware [RP21]. Quantifying [DRC^{+}18]. Quantitative [BCTV19, PPK18, ZJP21]. Queries [TBKJ19]. Query [YNL^{+}22]. Queue [THR20]. QuickLoc [TKP21].

Radio [GRA^{+}21]. Railway [CCTS20, LTTY18, LTT^{+}20]. Railways [ZMC^{+}23]. Ramp [FWL23]. Random [YFS^{+}19]. Range [YNL^{+}22]. Ransomware [KFH^{+}24]. RAP [TSC21]. Rate [PPK18]. Rate-Adaptive [PPK18]. RBM [WZZ^{+}20]. RBM-LSTM [WZZ^{+}20]. Reactive [BZS22]. Real [AAH^{+}19, BZW^{+}20, BXL^{+}18, GKEQ21, HMBP22, HSL^{+}19, LRN^{+}19, WGL19, WWW20, WLTW22, ZHZ19, ZMK^{+}17]. Real-Time [AAH^{+}19, BXL^{+}18, HSL^{+}19, LRN^{+}19, WGL19, WWW20, WLTW22, ZMK^{+}17, BZW^{+}20, GKEQ21, HMBP22]. Realization [MVS^{+}20, WZ17]. Reasoning [WJL^{+}20]. Reasoning-based [WJL^{+}20]. Recognition [CKP^{+}22, MLG24]. Reconfiguration [HKS17, MPP^{+}24]. Reduction [XCL^{+}19]. Reference [DHS^{+}24, DFK^{+}24]. References [BDL^{+}24]. Regions [AEA^{+}22]. regulating [AK20]. Regulation [MGH^{+}23]. Rehabilitation [BGM^{+}18]. Reinforcement [CLL^{+}21, FPK^{+}23, KMWB19, MMSM21, VLPB23, WCZ^{+}24]. Relay [LLZ^{+}21]. Relevance [ADD^{+}20]. Reliability [LLC^{+}19, PM19]. Reliable [FWL23, MYL^{+}20]. Remain [PMAM21]. Remote [CLW^{+}24, MLG24, ZMC^{+}23]. Renewable [HZI^{+}17]. Repair [BCTV19, MEM23]. Replay [CLW^{+}24, PMAM21]. Reproducible [PSM24]. Requirements [SDZV19]. Research [JHZ^{+}21, Sta17]. Residential [FBS^{+}21]. Resilience [KFH^{+}24, MPP^{+}24]. Resilient [AK20, ANC^{+}23, PDK^{+}18, RTY^{+}19, YZF18]. Resolution [YMH^{+}22]. Resource [LJP20, PMKD22, TSC21, XZA^{+}19]. Resource-Constrained [LJP20, TSC21]. Resource-Cost-Aware [XZA^{+}19]. Responsive [MMSM21]. Restart [JAR^{+}20]. Results [KFH^{+}24]. Retrofits [AEA^{+}22]. Reverse [BGGT21]. Review [CBB22]. Rich [DN23]. RIDS [GRA^{+}21]. Ring [CBB22, PSW^{+}21]. Ring-Based [PSW^{+}21]. Risk [AWB^{+}23, QXZ^{+}24, PLLH22]. Risk-Aware [QXZ^{+}24]. Road [AR19, CBB22, YS20]. Roadrunner [WWC22]. Robot [PRM^{+}20, YWM^{+}19]. Robot/Smart [PRM^{+}20]. Robotic [HSL^{+}19]. Robust [BBCV21, KMA^{+}22, MHP^{+}21, WHS18, ZJP21]. Robustness [KDM^{+}23]. RollBack [CLW^{+}24]. Rooms [JRGB^{+}18, WLTW22]. Route [TYT^{+}23]. Routing [GAT20]. Rover [HTU22]. Rover-copter [HTU22]. RSimplex [WHS18]. Runtime [HCT^{+}23, MVS^{+}20, SBW20].

Safe [EEC^{+}21, MGH^{+}23, WCZ^{+}24, WSH^{+}24, ZMC^{+}23, BZW^{+}20]. Safer [CHJ^{+}22]. Safety [AY17, BZW^{+}20, BHX^{+}20, CJSM23, CJSM24, DN23, FWL23, HMBP22, KDM^{+}23, ANC^{+}23, SCB^{+}20, SNBR23, TNY17, XBW^{+}19, XZA^{+}19]. Safety-critical [KDM^{+}23]. SafeWatch [BHX^{+}20]. Sampling [CGCH18, WKS18].

Sampling-Based [WKS18]. Scalable [ITB^{+}24, SDZV19]. Scale [SDZV19, ZHZ19]. Scales [BBS21]. Scaling [MBH^{+}22]. SchedGuard [CKT^{+}23]. Schedule [CKT^{+}23]. Scheduler [MEM23]. Schedules [IKG^{+}19]. Scheduling [GS18]. Scheme [HZW^{+}20, RPK^{+}21]. Scientific [Lu21]. Secrecy [SEB21, WWW20]. Secret [LNE^{+}20]. Section [AO22]. Secure [Asp20, HSI^{+}20, LLZ^{+}21, RSJ21, WJL^{+}20]. Securing [LYWY19, WLA19]. Security [AY17, ACLY21, AWB^{+}23, ASGB23, BBCV20, BTB^{+}18, CJSM23, CJSM24, DN23, HMBP22, KMM^{+}21, KDM^{+}23, LJP20, MJA^{+}23, PIW^{+}17, PYD^{+}20, SST^{+}23, SCB^{+}20, SNBR23, VBV^{+}19, VTK24, WKS18, WTZ^{+}19]. Security-Centric [WTZ^{+}19]. Selected [AO22]. Self [BLD^{+}20, CZSL20, EB20, JLD20, MLW20, MGH^{+}23, PXH^{+}19, ST18, SLJ^{+}23]. Self-Aware [JLD20, BLD^{+}20, CZSL20, EB20]. Self-awareness [MLW20]. Self-Driving [MGH^{+}23]. Self-powered [PXH^{+}19]. Self-Stabilizing [ST18]. Self-triggered [SLJ^{+}23]. Semantic [DFK^{+}24]. Semantically [DN23]. Semantics [BFV21, ZJP21]. sensed [APB^{+}21]. Sensing [BGGD20, MVS^{+}19, WGH^{+}17]. Sensing-based [BGGD20]. Sensor [KS22, LRN^{+}19, PIW^{+}17, SBW20, SLJ^{+}23, WD21]. Sensors [AR22]. Sequences [SNBR23]. Series [WZS^{+}19]. Service [LCK^{+}19, YS21, ZZZ^{+}17b]. Services [GAT20, WYW^{+}19, YMH^{+}22]. Set [YFS^{+}19]. Sets [MPA^{+}20]. Severity [KOTD21]. Shaping [LRN^{+}19]. Shared [GAT20]. Sharing [HZI^{+}17, RSJ21, WSH^{+}24, CLL^{+}21]. Shift [YKDL24]. Shifts [IIL^{+}17]. Short [GWO^{+}20]. Short-term [GWO^{+}20]. Side [TMF24]. Side-channel [TMF24]. Signal [BB21, HHS19, ZJP21]. Simplex [MGH^{+}23]. Simulation [JSY^{+}19, WEL^{+}22]. Simulink [BGGT21]. Simultaneous [MSS18]. Single [MM23]. Single-Vehicle [MM23]. Skip [WEL^{+}22]. SLATS [MSS18]. Slip [MH23]. Small [IIL^{+}17]. Smart [AR22, BAU20, CHJ^{+}22, CKP^{+}22, GRA^{+}21, HZI^{+}17, ITB^{+}24, IKG^{+}19, KOTD21, KRG17, LLC^{+}19, MPA^{+}20, PSW^{+}21, PRM^{+}20, PMKD22, SZCL23, TNY17, TYT^{+}23, TBKJ19, WZS^{+}19, YMH^{+}22]. Smart-Grid [BAU20]. Smartwatches [GVM^{+}20]. Smooth [MMSM21]. Social [LYWY19, SEB21, WYW^{+}19, YWZ^{+}19]. Socially [YWM^{+}19]. Software [GKEQ21, JAR^{+}20, KDM^{+}23, TSC21]. Solar [FBS^{+}21]. Solution [PYD^{+}20]. Soundness [LJZ18]. Source [ZZZ^{+}17a, ZHZ19]. Space [DYKE24, RRK^{+}22, BVS^{+}21]. Spaces [CHJ^{+}22, NVS^{+}21]. Spatio [MMTA24]. Spatio-Temporal [MMTA24]. Spatiotemporal [GWO^{+}20]. Special [AO22, AAH^{+}19, ACLY21, CH20, CJSM23, CJSM24, HZJ21, HZJ22, JLD20, LP18, MV24, SCB^{+}20, SD21, WZR^{+}19, ZA18a, ZA18b]. Specification [KMS20, MEM23, NHB^{+}18]. Specifications [HTU22]. Speed [MM23, YS20]. Spillback [WWC22]. Spot [LTT^{+}20]. Stability [BMJ^{+}20, MBH^{+}22]. Stabilizing [ST18]. Stadium [SSV^{+}18]. State [BVS^{+}21, MVS^{+}20, RP21]. State-space [BVS^{+}21]. Stateflow [JSY^{+}19]. Station [CCTS20]. Statistical [QXZ^{+}24, YKDL24]. Steal [PMAM21]. Steerable [FPK^{+}23]. Stochastic [LLC^{+}19, YZF18]. Storage [IIL^{+}17]. Store [KMM^{+}21]. Store-based [KMM^{+}21]. Strategies [GBH^{+}17]. Streaming [ARC^{+}23]. Streams [WMC18]. Strengthening [GBH^{+}17]. Structured [SBW20, WGH^{+}17]. Study [GGB^{+}18, HKS17, PKF24, PSM24, SJT^{+}22, TSC21]. Subscribe [ST18]. Suite [JRGB^{+}18]. Supercapacitor [CZSL20].

Supercapacitor-powered [CZSL20]. Supervision [DYKE24]. Support [KDM^{+}23, PRM^{+}20]. supported [BMS^{+}22]. Supporting [BTB^{+}18]. Surface [SST^{+}23]. Surrogate [QXZ^{+}24]. Surveillance [FPK^{+}23, YFS^{+}19]. Survey [AY17, JHZ^{+}21, KMA^{+}20, LSN^{+}18, MPA^{+}20, ROWA19, SZCL23]. Sustainable [LCC^{+}19]. Switching [GKMA22, YZF18]. Synchronization [MSS18, RTY^{+}19]. Synchronous [AR19, MM23]. Synergy [HMWZ18]. Syntactic [FGZ^{+}20]. Synthesis [BTB^{+}18, RZC^{+}22, YL23]. System [ADD^{+}20, BAU20, BHX^{+}20, BMS^{+}22, DAM^{+}18, FGZ^{+}20, GKEQ21, GRA^{+}21, HCT^{+}23, JAR^{+}20, JRGB^{+}18, KAR22, LCC^{+}19, LLZ^{+}21, LYA20, MSZ^{+}23, MPP^{+}24, NRB^{+}18, SSV^{+}18, WZZ^{+}20, YFS^{+}19, YAR^{+}18]. System-level [BAU20, GKEQ21]. Systematic [KMM^{+}21, SJT^{+}22]. Systematically [BXL^{+}18]. Systems [AAH^{+}19, ACLY21, ASGB23, BFV21, BBCV20, BH21, BVS^{+}21, BLD^{+}20, BDL^{+}24, BTF^{+}21, ByKLS19, BGGT21, BXL^{+}18, CZV^{+}19, CZSL20, CH20, CLL^{+}21, CLW^{+}24, DHS^{+}24, DFK^{+}24, ETM^{+}23, EB20, GP20, GBH^{+}17, HMBP22, HZJ21, HZJ22, HHS19, ITB^{+}24, JAR^{+}20, JLD20, JCO21, KS22, KOTD21, KMA^{+}22, KFH^{+}24, KSK20, LTTY18, Lee17, LP18, LJP20, LNE^{+}20, LTT^{+}20, LJZ18, LCK^{+}19, Lu21, MGL^{+}19, MLG24, ANC^{+}23, MHP^{+}21, MJA^{+}23, MRB^{+}22, MYL^{+}20, PLLH22, PMAM21, PKF24, PIW^{+}17, PDK^{+}18, RRK^{+}22, ROWA19, RP21, SJT^{+}22, SS22, SDZV19, Sta17, VTK24, WAH22, WLA19, WGH^{+}17, WMC18, WZS^{+}19, WZR^{+}19, WKS18, XCL^{+}19, XBW^{+}19, XZA^{+}19, YHL^{+}22, YKDL24, YZF18, YWZ^{+}19, YL23, ZYCL19, ZA18a, ZA18b, ZJP21, ZMK^{+}17, ZGAB21]. Systems-of-Systems [CZV^{+}19].

Taming [BVS^{+}21]. TangleCV [RSJ21].

Target [ZDY^{+}20]. Task [HSL^{+}19]. TCPS [SD21]. Technique [RSJ21]. Techniques [LSN^{+}18, VTK24]. Tele [BGM^{+}18]. Tele-Rehabilitation [BGM^{+}18]. Teleoperated [YHL^{+}22]. Temperatures [WLTW22]. Temporal [BB21, HTU22, MMTA24, ZJP21]. Tensor [WYW^{+}19]. Tensor-Train [WYW^{+}19]. Term [WMC18, GWO^{+}20]. Test [KMS20]. Testing [AGG^{+}21]. Theft [LZZD24]. Theory [BFV21, MYK^{+}22, PLLH22, YS21, YFS^{+}19]. Theory-Based [MYK^{+}22]. Therefore [EB20]. Thermostats [IKG^{+}19]. Things [WZ17, BTB^{+}18, DRC^{+}18, LSN^{+}18, PYD^{+}20, WZ17, ZA18a, ZA18b]. Think [EB20]. Threat [CZV^{+}19, THR20]. Threats [LTT^{+}20]. Time [AAH^{+}19, BXL^{+}18, CLW^{+}24, FWL23, GKEQ21, HFAD^{+}24, HSL^{+}19, KLMS20, LJZ18, LRN^{+}19, MSS18, PM19, WZS^{+}19, WGL19, WWW20, WLTW22, YS20, ZHZ19, ZMK^{+}17, BZW^{+}20, HMBP22, SD21]. Time-Agnostic [CLW^{+}24]. Time-aware [KLMS20]. Time-Constrained [PM19]. Time-Critical [LJZ18]. Time-Soundness [LJZ18]. Timed [KFSL19]. Timing [KMA^{+}22, KFSL19, ANC^{+}23, PSM24]. Timing-Based [ANC^{+}23, PSM24]. Token [PSW^{+}21]. Tolerance [JAR^{+}20]. Tolerant [TBKJ19, XZA^{+}19, ZLSZ20]. Tool [RZC^{+}22]. TOP [YS20]. Topology [LLC^{+}19]. TORUS [SDZV19]. TPrune [MYL^{+}21]. Traceability [SDZV19]. Traces [YBKJ22]. Tracking [BKIS19, BHX^{+}20, BGM^{+}18]. Traction [LTTY18]. Trade [STB^{+}18]. Trade-Based [STB^{+}18]. Tradeoff [DRC^{+}18]. Tradeoffs [HSL^{+}19]. trading [EEC^{+}21]. Traffic [BGGD20, CBB22, GP20, LRN^{+}19, MM23, NC21, THR20, VLPB23, WSH^{+}24, ZHZ19]. Traffic-type [GP20]. Train [MMSM21, WYW^{+}19]. Training [LZZD24]. Trajectories [YS20]. Transactions [CHJ^{+}22]. Transactive [EEC^{+}21]. Transceiver [BGMM19]. Transformer [MYL^{+}21]. Transient [PIW^{+}17]. Transit [ZZZ^{+}17b]. Transitions [MVS^{+}20]. Transmission [HZI^{+}17, LTT^{+}20]. Transmissions [TdAGI^{+}22]. Transportation [CH20, GWO^{+}20, ITB^{+}24, KSK20, NC21]. Travel [YS20]. Tree [MRL^{+}23]. Trees [BB21]. Trigger [NRB^{+}18]. Triggered [FCQY23, MH23, TdAGI^{+}22, SLJ^{+}23]. TruckSTM [MVS^{+}20]. True [WEL^{+}22]. Trust [APB^{+}21, LYWY19, MSZ^{+}23, SPH^{+}22]. Trustworthiness [SJT^{+}22]. Trustworthy [AAD^{+}21]. TSN [GP20]. TSN-based [GP20]. Tuning [MYK^{+}22]. Two [HMWZ18, MM23, TLW^{+}19]. Two-Level [TLW^{+}19]. Two-Speed [MM23]. type [GP20].

UACFinder [FGZ^{+}20]. UAS [RPK^{+}21]. UAV [KMWB19]. UAVs [MYK^{+}22]. Ultra [PXH^{+}19]. Ultra-low [PXH^{+}19]. Unacknowledged [PM19]. Unauthorized [GVM^{+}20]. Uncertain [HTU22]. Uncertainties [MVS^{+}19]. Understanding [ASGB23]. Undetected [PMAM21]. Unforeseen [MGH^{+}23]. Uniform [CGCH18]. Unpredictable [WSH^{+}24]. Unspecified [FGZ^{+}20]. Unsupervised [LZZD24, ZGAB21]. Update [HTU22]. Urban [ZZZ^{+}17a, ZZZ^{+}17b]. usage [AEA^{+}22]. User [AK20, APB^{+}21, HHS19, HZW^{+}20, LCC^{+}19, NVS^{+}21, PYD^{+}20, SCB^{+}20, WWW20]. User-Behavior-Aware [LCC^{+}19]. User-centeric [APB^{+}21]. User-Centric [SCB^{+}20, AK20, HZW^{+}20, PYD^{+}20]. Using [ARC^{+}23, BB21, CGCH18, CLL^{+}21, CKT^{+}23, GWO^{+}20, JCO21, LLZ^{+}21, RTY^{+}19, RRK^{+}22, SBW20, TSC21, YHL^{+}22, YFS^{+}19, AEA^{+}22, FPK^{+}23, QXZ^{+}24, ZMC^{+}23]. Utility [DRC^{+}18].

Utilizing [YS21].

V2V [BZW^{+}20]. VAE [RRK^{+}22]. Validating [BBS21]. Validation [SSV^{+}18, YBKJ22, YAR^{+}18, YL23]. VANETs [HSI^{+}20]. Vehicle [BCDT24, FCQY23, KS22, KMS20, MM23, MHP^{+}21, RKAG21, SBA21, VMDJ20, YS20]. Vehicles [AR19, GKMA22, KLMS20, KMA^{+}20, MSZ^{+}23, MGH^{+}23, MVS^{+}20, RSJ21, SPH^{+}22, VLPB23, WWC22, YS21]. Vehicular [Asp20, BGGD20, LNE^{+}20]. Velocity [MGH^{+}23]. Verifiable [PSW^{+}21]. Verification [Asp20, BZS22, ETM^{+}23, GGB^{+}18, HCT^{+}23, MEM23, PKF24, PPK18, QXZ^{+}24, SBA21]. Verified [JSY^{+}19]. via [DYKE24, GBH^{+}17, HKS17, JAR^{+}20, LHS^{+}21, MMSM21, WCZ^{+}24, WSH^{+}24]. Video [ARC^{+}23]. VIoLET [BBS21]. Virtual [BKIS19, KMS20]. Visual [AGG^{+}21, LCK^{+}19, VBV^{+}19]. Volatile [BGMM19, PXH^{+}19]. Voltage [RTY^{+}19].

Water [AEA^{+}22, PMAM21]. Water-usage [AEA^{+}22]. Watermarking [KSK20]. Watermarking-based [KSK20]. Wave [BGMM19]. Weak [DYKE24]. Weakly [HCT^{+}23]. Weakly-Hard [HCT^{+}23]. Wearable [BHX^{+}20, HMWZ18]. Wide [BTF^{+}21]. WiFi [XCL^{+}19]. Wind [AK20]. WInternet [WZ17]. Wired [LRN^{+}19]. Wireless [BMJ^{+}20, FWL23, GRA^{+}21, HBD^{+}20, MGL^{+}19, MLW20, MBH^{+}22, MYL^{+}20, SBW20, Sta17, TdAGI^{+}22, WLA19, WD21, YS21, ZMK^{+}17]. Worker [ZMC^{+}23]. working [NVS^{+}21]. WSNs [PM19, WWW20].

Yoneda [BGF22].

Z [BGMM19]. Z-Wave [BGMM19]. ZW0301 [BGMM19].

References

Ardagna:2021:TDT

[AAD^{+}21] Claudio A. Ardagna, Rasool Asal, Ernesto Damiani, Nabil El Ioini, Mehdi Elahi, and Claus Pahl. From trustworthy data to trustworthy IoT: a data collection methodology based on blockchain. ACM Transactions on Cyber-Physical Systems (TCPS), 5(1):11:1– 11:26, January 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3418686.

Almeida:2019:ISI

[AAH^{+}19] Luis Almeida, Bjorn Andersson, Jen-Wei Hsieh, Li-Pin Chang, and Xiaobo Sharon Hu. Introduction to the special issue on real-time aspects in cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 3(3):24:1– 24:2, October 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3342564.

Ambrosin:2021:ISI

[ACLY21] Moreno Ambrosin, Mauro Conti, Riccardo Lazzeretti, and Chia-Mu Yu. Introduction to the special issue on security and privacy for connected cyberphysical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 5(1):1:1–1:2, January 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3431201.

Alaeiyan:2020:MFR

[ADD^{+}20] Mohammadhadi Alaeiyan, Ali Dehghantanha, Tooska Dargahi, Mauro Conti, and Saeed Parsa. A multilabel fuzzy relevance clustering system for malware attack attribution in the edge layer of cyber-physical networks. ACM Transactions on Cyber-Physical Systems (TCPS), 4(3): 31:1–31:22, March 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3351881.

Abbas:2022:AEW

[AEA^{+}22] Samar Abbas, Ahmed Ehsan, Saad Ahmed, Sheraz Ali Khan, Tariq M. Jadoon, and Muhammad Hamad Alizai. ASHRAY: Enhancing water-usage comfort in developing regions using data-driven IoT retrofits. ACM Transactions on Cyber-Physical Systems (TCPS), 6(2): 18:1–18:??, April 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3491242.

Abrecht:2021:TDL

[AGG^{+}21] Stephanie Abrecht, Lydia Gauerhof, Christoph Gladisch, Konrad Groh, Christian Heinzemann, and Matthias Woehrle. Testing deep learning-based visual perception for automated driving.

ACM Transactions on Cyber-Physical Systems (TCPS), 5 (4):37:1–37:28, October 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3450356.

Arani:2020:UCA

[AK20] Mohammadreza Arani and Deepa Kundur. A usercentric approach toward resilient frequency-regulating wind generators. ACM Transactions on Cyber-Physical Systems (TCPS), 4(3):36:1–36:23, May 2020. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3389682.

Maruf:2023:TBF

[ANC^{+}23] Abdullah Al Maruf, Luyao Niu, Andrew Clark, J. Sukarno Mertoguno, and Radha Poovendran. A timing-based framework for designing resilient cyberphysical systems under safety constraint. ACM Transactions on Cyber-Physical Systems (TCPS), 7(3):19:1–19:??, July 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3594638.

AlFaruque:2022:ISS

[AO22] Mohammad Al Faruque and Meeko Mitsuko Oishi. Introduction to the special section on selected papers from ICCPS 2021. ACM Transactions on Cyber-Physical Systems (TCPS), 6(4):

29:1–29:??, October 2022. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3564157.

Azad:2021:PPC

[APB^{+}21] Muhammad Ajmal Azad, Charith Perera, Samiran Bag, Mahmoud Barhamgi, and Feng Hao. Privacy-preserving crowdsensed trust aggregation in the user-centeric Internet of People networks. ACM Transactions on Cyber-Physical Systems (TCPS), 5(1):4:1–4:24, January 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3390860.

Aoki:2019:CSP

[AR19] Shunsuke Aoki and Ragunathan (Raj) Rajkumar. CSIP: a synchronous protocol for automated vehicles at road intersections. ACM Transactions on Cyber-Physical Systems (TCPS), 3(3):25:1–25:25, October 2019. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/3226032.

Agarwal:2022:NAD

[AR22] Anshul Agarwal and Krithi Ramamritham. A novel approach for deploying minimum sensors in smart buildings. ACM Transactions on Cyber-Physical Systems (TCPS), 6(1): 2:1–2:29, January 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3477929.

Alsoliman:2023:IDF

[ARC^{+}23] Anas Alsoliman, Giulio Rigoni, Davide Callegaro, Marco Levorato, Cristina M. Pinotti, and Mauro Conti. Intrusion detection framework for invasive FPV drones using video streaming characteristics. ACM Transactions on Cyber-Physical Systems (TCPS), 7(2):12:1–12:??, April 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3579999.

Asiri:2023:UIC

[ASGB23] Mohammed Asiri, Neetesh Saxena, Rigel Gjomemo, and Pete Burnap. Understanding indicators of compromise against cyber-attacks in industrial control systems: a security perspective. ACM Transactions on Cyber-Physical Systems (TCPS), 7(2):15:1–15:??, April 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3587255.

Asplund:2020:CDV

[Asp20]  Mikael Asplund.  Combin-

ing detection and verification for secure vehicular cooperation groups. ACM Transactions on Cyber-Physical Systems (TCPS), 4(1):10:1– 10:31, January 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3322129.

Angermeier:2023:SRA

[AWB^{+}23] Daniel Angermeier, Hannah Wester, Kristian Beilke, Gerhard Hansch, and Jörn Eichler. Security risk assessments: Modeling and risk level propagation. ACM Transactions on Cyber-Physical Systems (TCPS), 7(1): 8:1–8:??, January 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3569458.

Altawy:2017:SPS

[AY17] Riham Altawy and Amr M. Youssef. Security, privacy, and safety aspects of civilian drones: A survey. ACM Transactions on Cyber-Physical Systems (TCPS), 1(2):7:1–7:25, February 2017. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL http:

//dl.acm.org/citation.cfm?
id=3001836.

Babun:2020:SLB

[BAU20] Leonardo Babun, Hidayet Aksu, and A. Selcuk Uluagac. A system-level behavioral detection framework for compromised CPS devices: Smart-grid case. ACM Transactions on Cyber-Physical Systems (TCPS), 4(2): 16:1–16:28, February 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3355300.

Bombara:2021:OOL

[BB21] Giuseppe Bombara and Calin Belta. Offline and online learning of signal temporal logic formulae using decision trees. ACM Transactions on Cyber-Physical Systems (TCPS), 5(3): 22:1–22:23, July 2021. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3433994.

Bakker:2020:HCP

[BBCV20] Craig Bakker, Arnab Bhattacharya, Samrat Chatterjee, and Draguna L. Vrabie. Hypergames and cyber-physical security for control systems. ACM Transactions on Cyber-Physical Systems (TCPS), 4(4):45:1– 45:41, August 2020. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3384676.

Bakker:2021:MHD

[BBCV21] Craig Bakker, Arnab Bhattacharya, Samrat Chatterjee, and Draguna L. Vrabie. Metagames and hypergames for deception-robust control. ACM Transactions on Cyber-Physical Systems (TCPS), 5(3): 23:1–23:25, July 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3439430.

Baheti:2021:VEE

[BBS21] Shrey Baheti, Shreyas Badiger, and Yogesh Simmhan. VIo-LET: an emulation environment for validating IoT deployments at large scales. ACM Transactions on Cyber-Physical Systems (TCPS), 5(3):25:1–25:39, July 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3446346.

Brighente:2024:EEV

[BCDT24] Alessandro Brighente, Mauro Conti, Denis Donadel, and Federico Turrin. EVScout2.0: Electric vehicle profiling through charging profile. ACM Transactions on Cyber-Physical Systems (TCPS), 8(2):11:1–11:??, April 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3565268.

Biagi:2019:MBQ

[BCTV19] Marco Biagi, Laura Carnevali, Fabio Tarani, and Enrico Vicario. Model-based quantitative evaluation of repair procedures in gas distribution networks. ACM Transactions on Cyber-Physical Systems (TCPS), 3(2): 19:1–19:26, March 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3284037.

Bengler:2024:RAH

[BDL^{+}24] Klaus Bengler, Werner Damm, Andreas Luedtke, Reiger Jochem, Benedikt Austel, Bianca Biebl, Martin Fränzle, Willem Hagemann, Moritz Held, David Hess, Klas Ihme, Severin Kacianka, Alyssa J. Kerscher, Laine Forrest, Sebastian Lehnhoff, Alexander Pretschner, Astrid Rakow, Daniel Sonntag, Janos Sztipanovits, Maike Schwammberger, Mark Schweda, Anirudh Unni, and Eric Veith. A references architecture for human cyber physical systems, Part II: Fundamental design principles for Human-CPS interaction. ACM Transactions on Cyber-Physical Systems (TCPS), 8(1): 3:1–3:??, January 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3622880.

Bakirtzis:2021:CSC

[BFV21] Georgios Bakirtzis, Cody H.

Fleming, and Christina Vasilakopoulou. Categorical semantics of cyber-physical systems theory. ACM Transactions on Cyber-Physical Systems (TCPS), 5(3):32:1–32:32, July 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3461669.

Bakirtzis:2022:YHA

[BGF22] Georgios Bakirtzis, Fabrizio Genovese, and Cody H. Fleming. Yoneda hacking: The algebra of attacker actions. ACM Transactions on Cyber-Physical Systems (TCPS), 6(3):25:1–25:??, July 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3531063.

Barnwal:2020:PDT

[BGGD20] Rajesh P. Barnwal, Nirnay Ghosh, Soumya K. Ghosh, and Sajal K. Das. Publish or drop traffic event alerts? Quality-aware decision making in participatory sensing-based vehicular CPS. ACM Transactions on Cyber-Physical Systems (TCPS), 4(1): 9:1–9:28, January 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3321480.

Bourbouh:2021:LSR

[BGGT21] Hamza Bourbouh, Pierre-Loïc Garoche, Christophe Garion, and Xavier Thirioux. From Lustre to Simulink: Reverse compilation for embedded systems applications. ACM Transactions on Cyber-Physical Systems (TCPS), 5(3):31:1–31:20, July 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3461668.

Buonocunto:2018:LTP

[BGM^{+}18] Pasquale Buonocunto, Andrea Giantomassi, Mauro Marinoni, Davide Calvaresi, and Giorgio Buttazzo. A limb tracking platform for tele-rehabilitation. ACM Transactions on Cyber-Physical Systems (TCPS), 2 (4):30:1–30:23, September 2018. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3148225.

Badenhop:2019:LUH

[BGMM19] C. W. Badenhop, S. R. Graham, B. E. Mullins, and L. O. Mailloux. Looking under the Hood of Z-Wave: Volatile memory introspection for the ZW0301 transceiver. ACM Transactions on Cyber-Physical Systems (TCPS), 3(2):20:1– 20:24, March 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3285030.

Barrere:2021:AMC

[BH21] Martín Barrère and Chris Hankin. Analysing mission-critical cyber-physical systems with AND/OR graphs and MaxSAT. ACM Transactions on Cyber-Physical Systems (TCPS), 5(3): 30:1–30:29, July 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3451169.

Bi:2020:SWH

[BHX^{+}20] Chongguang Bi, Jun Huang, Guoliang Xing, Landu Jiang, Xue Liu, and Minghua Chen. SafeWatch: a wearable hand motion tracking system for improving driving safety. ACM Transactions on Cyber-Physical Systems (TCPS), 4(1):13:1– 13:21, January 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3360323.

Barker:2019:BVP

[BKIS19] Sean Barker, Sandeep Kalra, David Irwin, and Prashant Shenoy. Building virtual power meters for online load tracking. ACM Transactions on Cyber-Physical Systems (TCPS), 3(2): 23:1–23:24, March 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3303860.

Bellman:2020:SAC

[BLD^{+}20] K. Bellman, C. Landauer, N. Dutt, L. Esterle, A. Herkersdorf, A. Jantsch, N. TaheriNejad, P. R. Lewis, M. Platzner, and K. Tammemäe. Self-aware cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 4(4):38:1– 38:26, August 2020. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3375716.

Baumann:2020:FFC

[BMJ^{+}20] Dominik Baumann, Fabian Mager, Romain Jacob, Lothar Thiele, Marco Zimmerling, and Sebastian Trimpe. Fast feedback control over multi-hop wireless networks with mode changes and stability guarantees. ACM Transactions on Cyber-Physical Systems (TCPS), 4(2): 18:1–18:32, February 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3361846.

Brik:2022:FSL

[BMS^{+}22] Bouziane Brik, Mourad Messaadia, M’hammed Sahnoun, Belgacem Bettayeb, and Mohamed Amin Benatia. Fogsupported low-latency monitoring of system disruptions in Industry 4.0: a federated learning approach. ACM Transactions on Cyber-Physical Systems (TCPS), 6(2):14:1–14:??, April 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3477272.

Banik:2023:AAL

[BRBB23] Sandeep Banik, Thiagarajan Ramachandran, Arnab Bhattacharya, and Shaunak D. Bopardikar. Automated adversaryin-the-loop cyber-physical defense planning. ACM Transactions on Cyber-Physical Systems (TCPS), 7(3):18:1–18:??, July 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3596222.

Bennaceur:2018:FDM

[BTB^{+}18] Amel Bennaceur, Thein Than Tun, Arosha K. Bandara, Yijun Yu, and Bashar Nuseibeh. Feature-driven mediator synthesis: Supporting collaborative security in the Internet of Things. ACM Transactions on Cyber-Physical Systems (TCPS), 2(3):

21:1–21:25, July 2018. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3134843.

Bhatia:2021:CCC

[BTF^{+}21] Laksh Bhatia, Ivana Tomić, Anqi Fu, Michael Breza, and Julie A. Mccann. Control communication co-design for wide area cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 5 (2):18:1–18:27, January 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3418528.

Bastos:2021:TSS

[BVS^{+}21] João Bastos, Jeroen Voeten, Sander Stuijk, Ramon Schiffelers, and Henk Corporaal. Taming the state-space explosion in the makespan optimization of flexible manufacturing systems. ACM Transactions on Cyber-Physical Systems (TCPS), 5(2): 15:1–15:26, January 2021. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3426194.

Bu:2018:SEC

[BXL^{+}18] Lei Bu, Wen Xiong, Chieh-Jan Mike Liang, Shi Han, Dongmei Zhang, Shan Lin, and Xuandong Li. Systematically ensuring the confidence of real-time home automation IoT systems. ACM Transactions on Cyber-Physical Systems (TCPS), 2(3): 22:1–22:23, July 2018. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3185501.

Bhuiyan:2019:DCP

[ByKLS19] Md Zakirul Alam Bhuiyan, Sy yen Kuo, Damian Lyons, and Zili Shao. Dependability in cyber-physical systems and applications. ACM Transactions on Cyber-Physical Systems (TCPS), 3(1):1:1– 1:4, January 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3271432.

Biggar:2022:MRC

[BZS22] Oliver Biggar, Mohammad Zamani, and Iman Shames. On modularity in reactive control architectures, with an application to formal verification. ACM Transactions on Cyber-Physical Systems (TCPS), 6(2): 19:1–19:??, April 2022. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3511606.

Bai:2020:MSM

[BZW^{+}20] Yunhao Bai, Kuangyu Zheng, Zejiang Wang, Xiaorui Wang, and Junmin Wang. MC-Safe: Multi-channel real-time V2V communication for enhancing driving safety. ACM Transactions on Cyber-Physical Systems (TCPS), 4(4):46:1– 46:27, August 2020. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3394961.

Chou:2022:LRR

[CBB22] Fang-Chieh Chou, Alben Rome Bagabaldo, and Alexandre M. Bayen. The lord of the ring road: a review and evaluation of autonomous control policies for traffic in a ring road. ACM Transactions on Cyber-Physical Systems (TCPS), 6(1): 8:1–8:25, January 2022. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3494577.

Cheh:2020:MAP

[CCTS20] Carmen Cheh, Binbin Chen, William G. Temple, and William H. Sanders. Modeling adversarial physical movement in a railway station: Classification and metrics. ACM Transactions on Cyber-Physical Systems (TCPS), 4(1):11:1– 11:25, January 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3349584.

Chang:2018:AAC

[CGCH18] Wanli Chang, Dip Goswami, Samarjit Chakraborty, and Arne Hamann. OS-aware automotive controller design using non-uniform sampling. ACM Transactions on Cyber-Physical Systems (TCPS), 2(4):26:1– 26:22, September 2018. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3121427.

Chakraborty:2020:ISI

[CH20] Samarjit Chakraborty and Tian He. Introduction to the special issue on transportation cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 4(1): 1:1–1:3, January 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3372495.

Chen:2022:ITS

[CHJ^{+}22] Chao Chen, Abdelsalam (Sumi)

Helal, Zhi Jin, Mingyue Zhang, and Choonhwa Lee. Io-Tranx: Transactions for safer smart spaces. ACM Transactions on Cyber-Physical Systems (TCPS), 6(1):1:1–1:26, January 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3471937.

Chakraborty:2023:ISI

[CJSM23] Samarjit Chakraborty, Somesh Jha, Soheil Samii, and Philipp Mundhenk. Introduction to the special issue on automotive CPS safety & security: Part 1. ACM Transactions on Cyber-Physical Systems (TCPS), 7(1): 1:1–1:??, January 2023. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3579986.

Chakraborty:2024:ISI

[CJSM24] Samarjit Chakraborty, Somesh Jha, Soheil Samii, and Philipp Mundhenk. Introduction to the special issue on automotive CPS safety & security: Part 2. ACM Transactions on Cyber-Physical Systems (TCPS), 8(2): 10:1–10:??, April 2024. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3650210.

Codispoti:2022:LNE

[CKP^{+}22] Jackson Codispoti, Atieh R. Khamesi, Nelson Penn, Simone Silvestri, and Eura Shin. Learning from non-experts: an interactive and adaptive learning approach for appliance recognition in smart homes. ACM Transactions on Cyber-Physical Systems (TCPS), 6(2):16:1–16:??, April 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3491241.

Chen:2023:SPA

[CKT^{+}23] Jiyang Chen, Tomasz Kloda, Rohan Tabish, Ayoosh Bansal, Chien-Ying Chen, Bo Liu, Sibin Mohan, Marco Caccamo, and Lui Sha. SchedGuard++: Protecting against schedule leaks using Linux containers on multicore processors. ACM Transactions on Cyber-Physical Systems (TCPS), 7(1):6:1–6:??, January 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3565974.

Chen:2021:DBD

[CLL^{+}21] Jianguo Chen, Kenli Li, Keqin Li, Philip S. Yu, and Zeng Zeng. Dynamic bicycle dispatching of dockless public bicyclesharing systems using multiobjective reinforcement learning. ACM Transactions on Cyber-Physical Systems (TCPS), 5(4): 34:1–34:24, October 2021. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3447623.

Csikor:2024:RNT

[CLW^{+}24] Levente Csikor, Hoon Wei Lim, Jun Wen Wong, Soundarya Ramesh, Rohini Poolat Parameswarath, and Mun Choon Chan. Roll-Back: a new time-agnostic replay attack against the automotive remote keyless entry systems. ACM Transactions on Cyber-Physical Systems (TCPS), 8(1):5:1–5:??, January 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3627827.

Chai:2020:SAP

[CZSL20] Ruizhi Chai, Ying Zhang, Geng Sun, and Hongsheng Li. Selfaware power management for maintaining event detection probability of supercapacitorpowered cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 4 (4):42:1–42:19, August 2020. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3375407.

Ceccarelli:2019:TAS

[CZV^{+}19] Andrea Ceccarelli, Tommaso Zoppi, Alexandr Vasenev, Marco Mori, Dan Ionita, Lorena Montoya, and Andrea Bondavalli. Threat analysis in systemsof-systems: an emergenceoriented approach. ACM Transactions on Cyber-Physical Systems (TCPS), 3(2):18:1– 18:24, March 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3234513.

DeVenuto:2018:MHS

[DAM^{+}18] D. De Venuto, V. F. Annese, G. Mezzina, F. Scioscia, M. Ruta, E. Di Sciascio, and A. Sangiovanni Vincentelli. A mobile health system for neurocognitive impairment evaluation based on P300 detection. ACM Transactions on Cyber-Physical Systems (TCPS), 2 (4):31:1–31:21, September 2018. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3140236.

Damm:2024:RAHb

[DFK^{+}24] Werner Damm, Martin Fränzle, Alyssa J. Kerscher, Forrest Laine, Klaus Bengler, Bianca Biebl, Willem Hagemann, Moritz Held, David Hess, Klas Ihme, Severin Kacianka, Sebastian Lehnhoff, Andreas Luedtke, Alexander Pretschner, Astrid Rakow, Jochem Rieger, Daniel Sonntag, Janos Sztipanovits, Maike Schwammberger, Mark Schweda, Alexander Trende, Anirudh Unni, and Eric Veith. A reference architecture of human cyber-physical systems — Part III: Semantic foundations. ACM Transactions on Cyber-Physical Systems (TCPS), 8(1): 4:1–4:??, January 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3622881.

Damm:2024:RAHa

[DHS^{+}24] Werner Damm, David Hess, Mark Schweda, Janos Sztipanovits, Klaus Bengler, Bianca Biebl, Martin Fränzle, Willem Hagemann, Moritz Held, Klas Ihme, Severin Kacianka, Alyssa J. Kerscher, Sebastian Lehnhoff, Andreas Luedtke, Alexander Pretschner, Astrid Rakow, Jochem Rieger, Daniel Sonntag, Maike Schwammberger, Benedikt Austel, Anirudh Unni, and Eric Veith. A reference architecture of human cyberphysical systems — Part I: Fundamental concepts. ACM Transactions on Cyber-Physical Systems (TCPS), 8(1):2:1–2:??, January 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3622879.

Dantas:2023:ASS

[DN23] Yuri Gil Dantas and Vivek Nigam. Automating safety and security co-design through semantically rich architecture patterns. ACM Transactions on Cyber-Physical Systems (TCPS), 7(1):5:1–5:??, January 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3565269.

Dong:2018:QUP

[DRC^{+}18] Roy Dong, Lillian J. Ratliff, Alvaro A. Cárdenas, Henrik Ohlsson, and S. Shankar Sastry. Quantifying the utility– privacy tradeoff in the Internet of Things. ACM Transactions on Cyber-Physical Systems (TCPS), 2(2):8:1– 8:28, June 2018. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3185511.

Das:2024:ILS

[DYKE24] Suman Das, Michael Yuhas, Rachel Koh, and Arvind Easwaran. Interpretable latent space for meteorological out-of-distribution detection via weak supervision. ACM Transactions on Cyber-Physical Systems (TCPS), 8(2): 24:1–24:??, April 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3651224.

Esterle:2020:TTY

[EB20] Lukas Esterle and John N. A.

Brown. I think therefore you are: Models for interaction in collectives of self-aware cyberphysical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 4(4):39:1–39:25, August 2020. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3375403.

Eisele:2021:SPF

[EEC^{+}21] Scott Eisele, Taha Eghtesad, Keegan Campanelli, Prakhar Agrawal, Aron Laszka, and Abhishek Dubey. Safe and private forward-trading platform for transactive microgrids. ACM Transactions on Cyber-Physical Systems (TCPS), 5(1): 8:1–8:29, January 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3403711.

Esiner:2023:MAP

[ETM^{+}23] Ertem Esiner, Utku Tefek, Daisuke Mashima, Binbin Chen, Zbigniew Kalbarczyk, and David M. Nicol. Message authentication and provenance verification for industrial control systems. ACM Transactions on Cyber-Physical Systems (TCPS), 7 (4):24:1–24:??, October 2023.

CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3607194.

Feng:2021:MDP

[FBS^{+}21] Menghong Feng, Noman Bashir, Prashant Shenoy, David Irwin, and Beka Kosanovic. Modeldriven per-panel solar anomaly detection for residential arrays. ACM Transactions on Cyber-Physical Systems (TCPS), 5(4): 40:1–40:20, October 2021. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3460236.

Fu:2023:PET

[FCQY23] Anqi Fu, Sijia Chen, Junfei Qiao, and Chengpu Yu. Periodic event-triggered CACC and communication co-design for vehicle platooning. ACM Transactions on Cyber-Physical Systems (TCPS), 7(4):28:1– 28:??, October 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3617125.

Fu:2020:UMS

[FGZ^{+}20] Zhicheng Fu, Chunhui Guo, Zhenyu Zhang, Shangping Ren, and Lui Sha. UACFinder: Mining syntactic carriers of unspecified assumptions in medical cyber-physical system design models. ACM Transactions on Cyber-Physical Systems (TCPS), 4(3):24:1– 24:25, March 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3375405.

Fahim:2023:ACS

[FPK^{+}23] Abdulrahman Fahim, Evangelos Papalexakis, Srikanth V. Krishnamurthy, Amit K. Roy Chowdhury, Lance Kaplan, and Tarek Abdelzaher. AcTrak: Controlling a steerable surveillance camera using reinforcement learning. ACM Transactions on Cyber-Physical Systems (TCPS), 7(2): 14:1–14:??, April 2023. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3585316.

Fan:2023:RWP

[FWL23] Xueli Fan, Qixin Wang, and Jie Liu. A reliable wireless protocol for highway and metered-ramp CAV collaborative merging with constanttime-headway safety guarantee. ACM Transactions on Cyber-Physical Systems (TCPS), 7 (4):23:1–23:??, October 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3609227.

Guan:2020:DRF

[GAT20] Yue Guan, Anuradha M. Annaswamy, and H. Eric Tseng. A dynamic routing framework for shared mobility services. ACM Transactions on Cyber-Physical Systems (TCPS), 4(1): 6:1–6:28, January 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3300181.

Gerostathopoulos:2017:SAC

[GBH^{+}17] Ilias Gerostathopoulos, Tomas Bures, Petr Hnetynka, Adam Hujecek, Frantisek Plasil, and Dominik Skoda. Strengthening adaptation in cyber-physical systems via meta-adaptation strategies. ACM Transactions on Cyber-Physical Systems (TCPS), 1(3):13:1–13:25, May 2017. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL http:

                                        //dl.acm.org/citation.cfm?
                                        id=2823345.

Geraldes:2018:FVM

[GGB^{+}18] André A. Geraldes, Luca Geretti, Davide Bresolin, Riccardo Muradore, Paolo Fiorini, Leonardo S. Mattos, and Tiziano Villa. Formal verification of medical CPS: a laser incision case study. ACM Transactions on Cyber-Physical Systems (TCPS), 2(4):35:1– 35:29, September 2018. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3140237.

Gemlau:2021:SLL

[GKEQ21] Kai-Björn Gemlau, Leonie KÖHLER, Rolf Ernst, and Sophie Quinton. System-level logical execution time: Augmenting the logical execution time paradigm for distributed real-time automotive software. ACM Transactions on Cyber-Physical Systems (TCPS), 5 (2):14:1–14:27, January 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3381847.

Ghafouri:2022:CCD

[GKMA22] Mohsen Ghafouri, Ekram Kabir, Bassam Moussa, and Chadi Assi. Coordinated charging and discharging of electric vehicles: a new class of switching attacks. ACM Transactions on Cyber-Physical Systems (TCPS), 6(3): 23:1–23:??, July 2022. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3524454.

Gavrilut:2020:TTA

[GP20] Voica Gavrilut and Paul Pop.

Traffic-type assignment for TSNbased mixed-criticality cyberphysical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 4(2):23:1–23:27, February 2020. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3371708.

Gimenez:2021:RRI

[GRA^{+}21] Pierre-François Gimenez, Jonathan Roux, Eric Alata, Guillaume Auriol, Mohamed Kaaniche, and Vincent Nicomette. RIDS: Radio intrusion detection and diagnosis system for wireless communications in smart environment. ACM Transactions on Cyber-Physical Systems (TCPS), 5 (3):24:1, July 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3441458.

Ghosh:2018:DSE

[GS18] Rajrup Ghosh and Yogesh Simmhan. Distributed scheduling of event analytics across edge and cloud. ACM Transactions on Cyber-Physical Systems (TCPS), 2(4):24:1–24:28, September 2018. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3140256.

Guerar:2020:CNA

[GVM^{+}20] Meriem Guerar, Luca Verderame, Alessio Merlo, Francesco Palmieri, Mauro Migliardi, and Luca Vallerini. CirclePIN: a novel authentication mechanism for smartwatches to prevent unauthorized access to IoT devices. ACM Transactions on Cyber-Physical Systems (TCPS), 4(3): 34:1–34:19, March 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3365995.

Gilanifar:2020:BSG

[GWO^{+}20] Mostafa Gilanifar, Hui Wang, Eren Erman Ozguven, Yuxun Zhou, and Reza Arghandeh. Bayesian spatiotemporal Gaussian process for short-term load forecasting using combined transportation and electricity data. ACM Transactions on Cyber-Physical Systems (TCPS), 4(1):2:1–2:25, January 2020. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/abs/10.1145/3300185.

Harbin:2020:APM

[HBD^{+}20] J. Harbin, A. Burns, R. I.

Davis, L. S. Indrusiak, I. Bate, and D. Griffin. The Air-Tight protocol for mixed criticality wireless CPS. ACM Transactions on Cyber-Physical Systems (TCPS), 4(2):19:1– 19:28, February 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3362987.

Hsieh:2023:SVR

[HCT^{+}23] Yi-Ting Hsieh, Tzu-Tao Chang, Chen-Jun Tsai, Shih-Lun Wu, Ching-Yuan Bai, Kai-Chieh Chang, Chung-Wei Lin, Eunsuk Kang, Chao Huang, and Qi Zhu. System verification and runtime monitoring with multiple weakly-hard constraints. ACM Transactions on Cyber-Physical Systems (TCPS), 7(3): 21:1–21:??, July 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3603380.

Hsueh:2024:DTD

[HFAD^{+}24] Kuei-Fang Hsueh, Ayleen Farnood, Isam Al-Darabsah, Mohammad Al Saaideh, Mohammad Al Janaideh, and Deepa Kundur. A deep time delay filter for cooperative adaptive cruise control. ACM Transactions on Cyber-Physical Systems (TCPS), 8(2): 17:1–17:??, April 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3631613.

Huang:2019:EEE

[HHS19] Hui Huang, Shiyan Hu, and Ye Sun. Energy-efficient ECG signal compression for user data input in cyber-physical systems by leveraging empirical mode decomposition. ACM Transactions on Cyber-Physical Systems (TCPS), 3(4):40:1– 40:19, October 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3341559.

He:2017:CSI

[HKS17] Liang He, Eugene Kim, and Kang G. Shin. A case study on improving capacity delivery of battery packs via reconfiguration. ACM Transactions on Cyber-Physical Systems (TCPS), 1(2):11:1–11:23, February 2017. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL http:

//dl.acm.org/citation.cfm?
id=3035539.

Hasan:2022:BJS

[HMBP22] Monowar Hasan, Sibin Mohan, Rakesh B. Bobba, and Rodolfo Pellizzoni. Beyond just safety: Delay-aware security monitoring for real-time control systems. ACM Transactions on Cyber-Physical Systems (TCPS), 6(3): 22:1–22:??, July 2022. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3520136.

Huang:2018:TBF

[HMWZ18] Qianyi Huang, Yan Mei, Wei Wang, and Qian Zhang. Toward battery-free wearable devices: The synergy between two feet. ACM Transactions on Cyber-Physical Systems (TCPS), 2(3): 20:1–20:18, July 2018. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3185503.

Hatzivasilis:2020:MSK

[HSI^{+}20] George Hatzivasilis, Othonas Soultatos, Sotiris Ioannidis, George Spanoudakis, Vasilios Katos, and Giorgos Demetriou. MobileTrust: Secure knowledge integration in VANETs. ACM Transactions on Cyber-Physical Systems (TCPS), 4(3): 33:1–33:25, March 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3364181.

Huang:2019:TNL

[HSL^{+}19] Pei-Chi Huang, Luis Sentis, Joel Lehman, Chien-Liang Fok, Aloysius K. Mok, and Risto Miikkulainen. Tradeoffs in neuroevolutionary learning-based real-time robotic task design in the imprecise computation framework. ACM Transactions on Cyber-Physical Systems (TCPS), 3(2): 14:1–14:29, March 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3178903.

Hashimoto:2022:CRC

[HTU22] Kazumune Hashimoto, Natsuko Tsumagari, and Toshimitsu Ushio. Collaborative Rovercopter path planning and exploration with temporal logic specifications based on Bayesian update under uncertain environments. ACM Transactions on Cyber-Physical Systems (TCPS), 6(2):11:1–11:??, April 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3470453.

Huang:2017:MTL

[HZI^{+}17] Zhichuan Huang, Ting Zhu, David Irwin, Aditya Mishra, Daniel Menasche, and Prashant Shenoy. Minimizing transmission loss in smart microgrids by sharing renewable energy. ACM Transactions on Cyber-Physical Systems (TCPS), 1(2):5:1–5:22, February 2017. CODEN ????

ISSN 2378-962X (print), 2378-9638 (electronic). URL http:

//dl.acm.org/citation.cfm?
id=2823355.

Hu:2021:ISI

[HZJ21] Jingtong Hu, Qi Zhu, and Susmit Jha. Introduction to the special issue on artificial intelligence and cyber-physical systems: Part 1. ACM Transactions on Cyber-Physical Systems (TCPS), 5(4):33:1–33:3, October 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3471164.

Hu:2022:ISI

[HZJ22] Jingtong Hu, Qi Zhu, and Susmit Jha. Introduction to the special issue on artificial intelligence and cyber-physical systems — Part 2. ACM Transactions on Cyber-Physical Systems (TCPS), 6(2):10:1–10:??, April 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3517045.

Huang:2020:HHU

[HZW^{+}20] Ke Huang, Xiaosong Zhang, Xiaofen Wang, Yi Mu, Fatemeh Rezaeibagha, Guangquan Xu, Hao Wang, Xi Zheng, Guomin Yang, Qi Xia, and Xiaojiang Du. HUCDO: a hybrid usercentric data outsourcing scheme. ACM Transactions on Cyber-Physical Systems (TCPS), 4(3): 35:1–35:23, May 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3379464.

Irwin:2017:EDE

[IIL^{+}17] David Irwin, Srinivasan Iyengar, Stephen Lee, Aditya Mishra, Prashant Shenoy, and Ye Xu. Enabling distributed energy storage by incentivizing small load shifts. ACM Transactions on Cyber-Physical Systems (TCPS), 1(2):10:1–10:30, February 2017. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL http:

                                        //dl.acm.org/citation.cfm?
                                        id=3015663.

Iyengar:2019:ISS

[IKG^{+}19] Srinivasan Iyengar, Sandeep Kalra, Anushree Ghosh, David Irwin, Prashant Shenoy, and Benjamin Marlin. Inferring smart schedules for dumb thermostats. ACM Transactions on Cyber-Physical Systems (TCPS), 3(2):17:1–17:29, March 2019. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/abs/10.1145/3226031.

Islam:2024:SPM

[ITB^{+}24] Md. Jaminur Islam, Jose Paolo Talusan, Shameek Bhattacharjee, Francis Tiausas, Abhishek Dubey, Keiichi Yasumoto, and Sajal K. Das. Scalable Pythagorean mean-based incident detection in smart transportation systems. ACM Transactions on Cyber-Physical Systems (TCPS), 8(2):20:1–20:??, April 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3603381.

Jagtap:2020:SFT

[JAR^{+}20] Pushpak Jagtap, Fardin Abdi, Matthias Rungger, Majid Zamani, and Marco Caccamo. Software fault tolerance for cyberphysical systems via full system restart. ACM Transactions on Cyber-Physical Systems (TCPS), 4(4):47:1–47:20, August 2020. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3407183.

Jesus:2021:UML

[JCO21] Gonçalo Jesus, António Casimiro, and Anabela Oliveira. Using machine learning for dependable outlier detection in environmental monitoring systems. ACM Transactions on Cyber-Physical Systems (TCPS), 5(3): 29:1–29:30, July 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3445812.

Jiang:2021:RPC

[JHZ^{+}21] Wei Jiang, Zhiyuan He, Jinyu Zhan, Weijia Pan, and Deepak Adhikari. Research progress and challenges on application-driven adversarial examples: a survey. ACM Transactions on Cyber-Physical Systems (TCPS), 5(4):

39:1–39:25, October 2021. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3470493.

Jantsch:2020:ISI

[JLD20] Axel Jantsch, Peter R. Lewis, and Nikil Dutt. Introduction to the special issue on selfaware cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 4(4): 37:1–37:2, August 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3397266.

Joerger:2018:CPS

[JRGB^{+}18] Guillaume Joerger, Juliette Rambourg, Helene Gaspard-Boulinc, Stephane Conversy, Barbara L. Bass, Brian J. Dunkin, and Marc Garbey. A cyber-physical system to improve the management of a large suite of operating rooms. ACM Transactions on Cyber-Physical Systems (TCPS), 2 (4):34:1–34:24, September 2018. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3140234.

Jiang:2019:DMD

[JSY^{+}19] Yu Jiang, Houbing Song, Yixiao Yang, Han Liu, Ming Gu, Yong Guan, Jiaguang Sun, and Lui Sha. Dependable model-driven development of CPS: From stateflow simulation to verified implementation. ACM Transactions on Cyber-Physical Systems (TCPS), 3(1): 12:1–12:31, January 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3078623.

Jiang:2017:EBM

[JZL^{+}17] Jian-Min Jiang, Huibiao Zhu, Qin Li, Yongxin Zhao, Lin Zhao, Shi Zhang, Ping Gong, Zhong Hong, and Donghuo Chen. Event-based mobility modeling and analysis. ACM Transactions on Cyber-Physical Systems (TCPS), 1(2):9:1–9:32, February 2017. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL http:

//dl.acm.org/citation.cfm?
id=2823353.

Khodabandehloo:2022:FCP

[KAR22] Elham Khodabandehloo, Abbas Alimohammadi, and Daniele Riboni. FreeSia: a cyber-physical system for cognitive assessment through frequency-domain indoor locomotion analysis. ACM Transactions on Cyber-Physical Systems (TCPS), 6(2):12:1– 12:??, April 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3470454.

Koley:2023:CSS

[KDM^{+}23] Ipsita Koley, Soumyajit Dey, Debdeep Mukhopadhyay, Sachin Singh, Lavanya Lokesh, and Shantaram Vishwanath Ghotgalkar. CAD support for security and robustness analysis of safety-critical automotive software. ACM Transactions on Cyber-Physical Systems (TCPS), 7(1):7:1–7:??, January 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3571287.

Kirkman:2024:BAI

[KFH^{+}24] Stephen S. Kirkman, Steven Fulton, Jeffrey Hemmes, Christopher Garcia, and Justin C. Wilson. A blockchain architecture to increase the resilience of industrial control systems from the effects of a ransomware attack: a proposal and initial results. ACM Transactions on Cyber-Physical Systems (TCPS), 8(1):9:1–9:??, January 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3637553.

Kim:2019:DTP

[KFSL19] Baekgyu Kim, Lu Feng, Oleg Sokolsky, and Insup Lee. Determining timing parameters for the code generation from platform-independent timed models. ACM Transactions on Cyber-Physical Systems (TCPS), 3(3):28:1–28:32, October 2019. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3230711.

Khayatian:2020:CTA

[KLMS20] Mohammad Khayatian, Yingyan Lou, Mohammadreza Mehrabian, and Aviral Shirvastava. Crossroads+: a timeaware approach for intersection management of connected autonomous vehicles. ACM Transactions on Cyber-Physical Systems (TCPS), 4(2):20:1– 20:28, February 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3364182.

Khayatian:2020:SIM

[KMA^{+}20] Mohammad Khayatian, Mohammadreza Mehrabian, Edward Andert, Rachel Dedinsky, Sarthake Choudhary, Yingyan Lou, and Aviral Shirvastava. A survey on intersection management of connected autonomous vehicles. ACM Transactions on Cyber-Physical Systems (TCPS), 4(4):48:1–48:27, August 2020. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3407903.

Khayatian:2022:PBD

[KMA^{+}22] Mohammad Khayatian, Mohammadreza Mehrabian, Edward Andert, Reese Grimsley, Kyle Liang, Yi Hu, Ian McCormack, Carlee Joe-Wong, Jonathan Aldrich, Bob Iannucci, and Aviral Shrivastava. Plan B: Design methodology for cyberphysical systems robust to timing failures. ACM Transactions on Cyber-Physical Systems (TCPS), 6(3):21:1–21:??, July 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3516449.

Kafle:2021:SCD

[KMM^{+}21] Kaushal Kafle, Kevin Moran, Sunil Manandhar, Adwait Nadkarni, and Denys Poshyvanyk. Security in centralized data store-based home automation platforms: a systematic analysis of Nest and Hue. ACM Transactions on Cyber-Physical Systems (TCPS), 5(1):2:1–2:27, January 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3418286.

Kim:2020:TSG

[KMS20] Baekgyu Kim, Takato Masuda, and Shinichi Shiraishi. Test specification and generation for connected and autonomous vehicle in virtual environments. ACM Transactions on Cyber-Physical Systems (TCPS), 4(1): 8:1–8:26, January 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3311954.

Koch:2019:RLU

[KMWB19] William Koch, Renato Mancuso, Richard West, and Azer Bestavros. Reinforcement learning for UAV attitude control. ACM Transactions on Cyber-Physical Systems (TCPS), 3(2): 22:1–22:21, March 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3301273.

Khaled:2021:ASS

[KOTD21] Abdelaziz Khaled, Samir Ouchani, Zahir Tari, and Khalil Drira. Assessing the severity of smart attacks in industrial cyberphysical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 5(1):10:1–10:28, January 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3422369.

Khandeparkar:2017:QDD

[KRG17] Kedar Khandeparkar, Krithi Ramamritham, and Rajeev Gupta. QoS-driven data processing algorithms for smart electric grids. ACM Transactions on Cyber-Physical Systems (TCPS), 1(3):14:1–14:24, May 2017. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL http:

//dl.acm.org/citation.cfm?
id=3047410.

Kang:2022:DMS

[KS22] Liuwang Kang and Haiying Shen. Detection and mitigation of sensor and CAN bus attacks in vehicle anti-lock braking systems. ACM Transactions on Cyber-Physical Systems (TCPS), 6(1):9:1–9:24, January 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3495534.

Ko:2020:DWB

[KSK20] Woo-Hyun Ko, Bharadwaj Satchidanandan, and P. R. Kumar. Dynamic watermarkingbased defense of transportation cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 4(1): 12:1–12:21, January 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3361700.

Kuo:2017:I

[Kuo17] Tei Kuo. Introduction. ACM Transactions on Cyber-Physical Systems (TCPS), 1(1):1e:1, February 2017. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL http:

                                        //dl.acm.org/citation.cfm?
                                        id=3047402.

Li:2019:SUB

[LCC^{+}19] Wei Li, Xiaomin Chang, Junwei Cao, Ting Yang, Yaojie Sun, and Albert Y. Zomaya. A sustainable and user-behavior-aware cyber-physical system for home energy management. ACM Transactions on Cyber-Physical Systems (TCPS), 3(4): 37:1–37:24, October 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3320075.

Liu:2019:DVL

[LCK^{+}19] Yang Liu, Xiaoming Chen, Dileep Kadambi, Ajinkya Bari, Xin Li, Shiyan Hu, and Pingqiang Zhou. Dependable visual light-based indoor localization with automatic anomaly detection for location-based service of mobile cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 3(1): 5:1–5:17, January 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3162051.

Lee:2017:FLC

[Lee17] Edward A. Lee. Fundamental limits of cyber-physical systems modeling. ACM Transactions on Cyber-Physical Systems (TCPS), 1(1):3:1–3:26, February 2017. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL http:

//dl.acm.org/citation.cfm?
id=2912149.

Laskar:2021:EIF

[LHS^{+}21] Md Tahmid Rahman Laskar, Jimmy Xiangji Huang, Vladan Smetana, Chris Stewart, Kees Pouw, Aijun An, Stephen Chan, and Lei Liu. Extending isolation forest for anomaly detection in big data via K-means. ACM Transactions on Cyber-Physical Systems (TCPS), 5(4): 41:1–41:26, October 2021. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3460976.

Lesi:2020:ISR

[LJP20] Vuk Lesi, Ilija Jovanov, and Miroslav Pajic. Integrating security in resource-constrained cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 4(3):28:1– 28:27, May 2020. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3380866.

Liu:2018:TST

[LJZ18] Guanjun Liu, Changjun Jiang, and Mengchu Zhou. Timesoundness of time Petri nets modelling time-critical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 2(2): 11:1–11:27, June 2018. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3185502.

Li:2019:RAS

[LLC^{+}19] Beibei Li, Rongxing Lu, Kim-Kwang Raymond Choo, Wei Wang, and Sheng Luo. On reliability analysis of smart grids under topology attacks: a stochastic Petri net approach. ACM Transactions on Cyber-Physical Systems (TCPS), 3(1): 10:1–10:25, January 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3127021.

Li:2021:BSA

[LLZ^{+}21] Kai Li, Ning Lu, Jingjing Zheng, Pei Zhang, Wei Ni, and Eduardo Tovar. BloothAir: a secure aerial relay system using Bluetooth connected autonomous drones. ACM Transactions on Cyber-Physical Systems (TCPS), 5(3): 27:1–27:22, July 2021. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3448254.

Li:2020:DIS

[LNE^{+}20] Kai Li, Wei Ni, Yousef Emami, Yiran Shen, Ricardo Severino, David Pereira, and Eduardo Tovar. Design and implementation of secret key agreement for platoon-based vehicular cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 4(2): 22:1–22:20, February 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3365996.

Lee:2018:GES

[LP18] Insup Lee and Miroslav Pajic.

Guest editorial: Special issue on medical cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 2(4): 29:1–29:2, September 2018. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3243122.

Loureiro:2019:EAR

[LRN^{+}19] João Loureiro, Raghuraman Rangarajan, Borislav Nikolic, Leandro Soares Indrusiak, and Eduardo Tovar. Extensive analysis of a real-time dense wired sensor network based on traffic shaping. ACM Transactions on Cyber-Physical Systems (TCPS), 3(3):27:1– 27:27, October 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3230872.

Liu:2018:SMC

[LSN^{+}18] Jinwei Liu, Haiying Shen, Husnu S. Narman, Wingyan Chung, and Zongfang Lin. A survey of mobile crowdsensing techniques: a critical component for the Internet of Things. ACM Transactions on Cyber-Physical Systems (TCPS), 2(3): 18:1–18:26, July 2018. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3185504.

Lim:2020:DIT

[LTT^{+}20] Hoon Wei Lim, William G. Temple, Bao Anh N. Tran, Binbin Chen, Zbigniew Kalbarczyk, and Jianying Zhou. Data integrity threats and countermeasures in railway spot transmission systems. ACM Transactions on Cyber-Physical Systems (TCPS), 4(1):7:1– 7:26, January 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3300179.

Lakshminarayana:2018:MDF

[LTTY18] Subhash Lakshminarayana, Teo Zhan Teng, Rui Tan, and David K. Y. Yau. Modeling and detecting false data injection attacks against railway traction power systems. ACM Transactions on Cyber-Physical Systems (TCPS), 2(4):28:1– 28:29, September 2018. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3226030.

Lu:2021:TSE

[Lu21] Chenyang Lu. Toward a scientific and engineering discipline of cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 5(3): 22e:1–22e:2, July 2021. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3464945.

Lu:2020:PBI

[LYA20] Hong Lu, Tao Yue, and Shaukat Ali. Pattern-based interactive configuration derivation for cyber-physical system product lines. ACM Transactions on Cyber-Physical Systems (TCPS), 4(4):44:1–44:24, August 2020. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3389397.

Li:2019:SCD

[LYWY19] Ning Li, Zheng Yan, Mingjun Wang, and Laurence T. Yang. Securing communication data in pervasive social networking based on trust with KP-ABE. ACM Transactions on Cyber-Physical Systems (TCPS), 3(1): 9:1–9:23, January 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3145624.

Liang:2024:UBB

[LZZD24] Qiushi Liang, Shengjie Zhao, Jiangfan Zhang, and Hao Deng. Unsupervised BLSTM-based electricity theft detection with training data contaminated. ACM Transactions on Cyber-Physical Systems (TCPS), 8(1): 1:1–1:??, January 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3604432.

Mager:2022:SBB

[MBH^{+}22] Fabian Mager, Dominik Baumann, Carsten Herrmann, Sebastian Trimpe, and Marco Zimmerling. Scaling beyond bandwidth limitations: Wireless control with stability guarantees under overload. ACM Transactions on Cyber-Physical Systems (TCPS), 6(3):20:1–20:??, July 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3502299.

Mousavi:2023:FSV

[MEM23] Hassan Mousavi, Ali Ebnenasir, and Elham Mahmoudzadeh. Formal specification, verification and repair of Contiki’s scheduler. ACM Transactions on Cyber-Physical Systems (TCPS), 7 (4):22:1–22:??, October 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3605948.

Mao:2023:SSS

[MGH^{+}23] Yanbing Mao, Yuliang Gu, Naira Hovakimyan, Lui Sha, and Petros Voulgaris. SL1-Simplex: Safe velocity regulation of selfdriving vehicles in dynamic and unforeseen environments. ACM Transactions on Cyber-Physical Systems (TCPS), 7(1): 2:1–2:??, January 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3564273.

Ma:2019:HCP

[MGL^{+}19] Yehan Ma, Dolvara Gunatilaka, Bo Li, Humberto Gonzalez, and Chenyang Lu. Holistic cyber-physical management for dependable wireless control systems. ACM Transactions on Cyber-Physical Systems (TCPS), 3(1):3:1– 3:25, January 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3185510.

Mamduhi:2023:ETC

[MH23] Mohammad H. Mamduhi and Ehsan Hashemi. Event-triggered control with intermittent communications over erasure channels for leader-follower problems with the combined-slip effect. ACM Transactions on Cyber-Physical Systems (TCPS), 7(4): 29:1–29:??, October 2023. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3625562.

Miao:2021:DDD

[MHP^{+}21] Fei Miao, Sihong He, Lynn Pepin, Shuo Han, Abdeltawab Hendawi, Mohamed E. Khalefa, John A. Stankovic, and George Pappas. Data-driven distributionally robust optimization for vehicle balancing of mobilityon-demand systems. ACM Transactions on Cyber-Physical Systems (TCPS), 5(2):17:1– 17:27, January 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3418287.

Mohamad:2023:CAD

[MJA^{+}23] Mazen Mohamad, Rodi Jolak, Örjan Askerdal, Jan-Philipp Steghöfer, and Riccardo Scandariato. CASCADE: an assetdriven approach to build security assurance cases for automotive systems. ACM Transactions on Cyber-Physical Systems (TCPS), 7(1):3:1–3:??, January 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3569459.

Man:2024:RPA

[MLG24] Yanmao Man, Ming Li, and Ryan Gerdes. Remote perception attacks against camerabased object recognition systems and countermeasures. ACM Transactions on Cyber-Physical Systems (TCPS), 8(2):14:1– 14:??, April 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3596221.

Ma:2020:EHC

[MLW20] Yehan Ma, Chenyang Lu, and Yebin Wang. Efficient holistic control: Self-awareness across controllers and wireless networks. ACM Transactions on Cyber-Physical Systems (TCPS), 4(4):41:1–41:27, August 2020. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3371500.

Murthy:2022:CPA

[MM22] Dharshan Krishna Murthy and Alejandro Masrur. A cyberphysical approach for emergency braking in close-distance driving arrangements. ACM Transactions on Cyber-Physical Systems (TCPS), 6(3):24:1–24:??, July 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3526117.

Markert:2023:TSS

[MM23] Daniel Markert and Alejandro Masrur. A two-speed synchronous traffic protocol for intelligent intersections: From single-vehicle to platoon crossing. ACM Transactions on Cyber-Physical Systems (TCPS), 7(2):10:1–10:??, April 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3571289.

Mysore:2021:HTY

[MMSM21] Siddharth Mysore, Bassel Mabsout, Kate Saenko, and Renato Mancuso. How to train your quadrotor: a framework for consistently smooth and responsive flight control via reinforcement learning. ACM Transactions on Cyber-Physical Systems (TCPS), 5(4):36:1– 36:24, October 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3466618.

Mortlock:2024:CCA

[MMTA24] Trier Mortlock, Arnav Malawade, Kohei Tsujio, and Mohammad Al Faruque. CASTNet: a context-aware, spatio-temporal dynamic motion prediction ensemble for autonomous driving. ACM Transactions on Cyber-Physical Systems (TCPS), 8(2): 23:1–23:??, April 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3648622.

Mai:2021:DLP

[MN21] Tieu Long Mai and Nicolas Navet. Deep learning to predict the feasibility of prioritybased Ethernet network configurations. ACM Transactions on Cyber-Physical Systems (TCPS), 5(4):45:1–45:26, October 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3468890.

Ma:2020:DSM

[MPA^{+}20] Meiyi Ma, Sarah M. Preum, Mohsin Y. Ahmed, William Tärneberg, Abdeltawab Hendawi, and John A. Stankovic. Data sets, modeling, and decision making in smart cities: a survey. ACM Transactions on Cyber-Physical Systems (TCPS), 4(2):14:1–14:28, February 2020. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3355283.

Matei:2024:SRT

[MPP^{+}24] Ion Matei, Wiktor Piotrowski, Alexandre Perez, Johan de Kleer, Jorge Tierno, Wendy Mungovan, and Vance Turnewitsch. System resilience through health monitoring and reconfiguration. ACM Transactions on Cyber-Physical Systems (TCPS), 8(1):

7:1–7:??, January 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3631612.

Mohammed:2022:CCO

[MRB^{+}22] Abubakar Sadiq Mohammed, Philipp Reinecke, Pete Burnap, Omer Rana, and Eirini Anthi. Cybersecurity challenges in the offshore oil and gas industry: an industrial cyber-physical systems (ICPS) perspective. ACM Transactions on Cyber-Physical Systems (TCPS), 6(3): 28:1–28:??, July 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3548691.

Mehta:2023:DDC

[MRL^{+}23] Jarul Mehta, Guillaume Richard, Loren Lugosch, Derek Yu, and Brett H. Meyer. DT-DS: CAN intrusion detection with decision tree ensembles. ACM Transactions on Cyber-Physical Systems (TCPS), 7(1):4:1–4:??, January 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3566132.

Martin:2018:SSL

[MSS18] Paul Martin, Andrew Symington, and Mani Srivastava. SLATS: Simultaneous localization and time synchronization. ACM Transactions on Cyber-Physical Systems (TCPS), 2(3): 19:1–19:25, July 2018. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3134841.

Mahmood:2023:TDT

[MSZ^{+}23] Adnan Mahmood, Quan Z.

Sheng, Wei Emma Zhang, Yan Wang, and Subhash Sagar. Toward a distributed trust management system for misbehavior detection in the Internet of vehicles. ACM Transactions on Cyber-Physical Systems (TCPS), 7(3):16:1–16:??, July 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3594637.

Mitra:2024:ISI

[MV24] Sayan Mitra and Nalini Venkatasubramanian. Introduction to special issue for ICCPS 2022. ACM Transactions on Cyber-Physical Systems (TCPS), 8(2): 18:1–18:??, April 2024. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3661449.

Medina:2019:DCI

[MVS^{+}19] Róbinson Medina, Juan Valencia, Sander Stuijk, Dip Goswami, and Twan Basten. Designing a controller with imagebased pipelined sensing and additive uncertainties. ACM Transactions on Cyber-Physical Systems (TCPS), 3(3):33:1– 33:26, October 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3326067.

Mukherjee:2020:TRR

[MVS^{+}20] Subhojeet Mukherjee, Jeffrey C. Van Etten, Namburi Rani Samyukta, Jacob Walker, Indrakshi Ray, and Indrajit Ray. TruckSTM: Runtime realization of operational state transitions for medium and heavy duty vehicles. ACM Transactions on Cyber-Physical Systems (TCPS), 4(1):4:1– 4:25, January 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3300183.

Moolchandani:2022:GTB

[MYK^{+}22] Diksha Moolchandani, Kishore Yadav, Geesara Kulathunga, Ilya Afanasyev, Anshul Kumar, Manuel Mazzara, and Smruti Sarangi. Game theory-based parameter tuning for energyefficient path planning on modern UAVs. ACM Transactions on Cyber-Physical Systems (TCPS), 6(4):34:1–34:??, October 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3565270.

Munir:2020:RCL

[MYL^{+}20] Sirajum Munir, Hao-Tsung Yang, Shan Lin, S. M. Shahriar Nirjon, Chen Lin, Enamul Hoque, John A. Stankovic, and Kamin Whitehouse. Reliable communication and latency bound generation in wireless cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 4(2):15:1– 15:26, February 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3354917.

Mao:2021:TET

[MYL^{+}21] Jiachen Mao, Huanrui Yang, Ang Li, Hai Li, and Yiran Chen. TPrune: Efficient transformer pruning for mobile devices. ACM Transactions on Cyber-Physical Systems (TCPS), 5(3): 26:1–26:22, July 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3446640.

Niu:2021:DPI

[NC21] Luyao Niu and Andrew Clark.

A differentially private incentive design for traffic offload to public transportation. ACM Transactions on Cyber-Physical Systems (TCPS), 5(2):20:1– 20:27, January 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3430847.

Nguyen:2018:CPS

[NHB^{+}18] Luan V. Nguyen, Khaza Anuarul Hoque, Stanley Bak, Steven Drager, and Taylor T. Johnson. Cyber-physical specification mismatches. ACM Transactions on Cyber-Physical Systems (TCPS), 2(4):23:1– 23:26, September 2018. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3170500.

Nacci:2018:BTA

[NRB^{+}18] Alessandro A. Nacci, Vincenzo Rana, Bharathan Balaji, Paola Spoletini, Rajesh Gupta, Donatella Sciuto, and Yuvraj Agarwal. BuildingRules: a trigger– action-based system to manage complex commercial buildings. ACM Transactions on Cyber-Physical Systems (TCPS), 2(2): 13:1–13:22, June 2018. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3185500.

Nagarathinam:2021:UPO

[NVS^{+}21] Srinarayana Nagarathinam, Arunchandar Vasan, Venkatesh Sarangan, Rajesh Jayaprakash, and Anand Sivasubramaniam. User placement and optimal cooling energy for co-working building spaces. ACM Transactions on Cyber-Physical Systems (TCPS), 5 (2):21:1–21:24, January 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3432818.

Pradhan:2018:CGD

[PDK^{+}18] Subhav Pradhan, Abhishek Dubey, Shweta Khare, Saideep Nannapaneni, Aniruddha Gokhale, Sankaran Mahadevan, Douglas C. Schmidt, and Martin Lehofer. CHARIOT: Goaldriven orchestration middleware for resilient IoT systems. ACM Transactions on Cyber-Physical Systems (TCPS), 2(3): 16:1–16:37, July 2018. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3134844.

Park:2017:SCP

[PIW^{+}17] Junkil Park, Radoslav Ivanov, James Weimer, Miroslav Pajic, Sang Hyuk Son, and Insup Lee. Security of cyber-physical systems in the presence of transient sensor faults. ACM Transactions on Cyber-Physical Systems (TCPS), 1(3):15:1–15:23, May 2017. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL http:

//dl.acm.org/citation.cfm?
id=3064809.

Panahi:2024:CPA

[PKF24] Vahid Panahi, Mehdi Kargahi, and Fathiyeh Faghih. Control performance analysis of automotive cyber-physical systems: a study on efficient formal verification. ACM Transactions on Cyber-Physical Systems (TCPS), 8(2):12:1–12:??, April 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3576046.

Pal:2022:HHC

[PLLH22] Ranjan Pal, Peihan Liu, Taoan Lu, and Ed Hua. How hard is cyber-risk management in IT/OT systems? A theory to classify and conquer hardness of insuring ICSs. ACM Transactions on Cyber-Physical Systems (TCPS), 6(4):35:1– 35:??, October 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3568399.

Parsch:2019:ARU

[PM19] Philip Parsch and Alejandro Masrur. Accounting for reliability in unacknowledged timeconstrained WSNs. ACM Transactions on Cyber-Physical Systems (TCPS), 3(3):26:1– 26:28, October 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3241746.

Palleti:2021:CRA

[PMAM21] Venkata Reddy Palleti, Vishrut Kumar Mishra, Chuadhry Mujeeb Ahmed, and Aditya Mathur. Can replay attacks designed to steal water from water distribution systems remain undetected? ACM Transactions on Cyber-Physical Systems (TCPS), 5(1): 9:1–9:19, January 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3406764.

Pettet:2022:HPD

[PMKD22] Geoffrey Pettet, Ayan Mukhopadhyay, Mykel J. Kochenderfer, and Abhishek Dubey. Hierarchical planning for dynamic resource allocation in smart and connected communities. ACM Transactions on Cyber-Physical Systems (TCPS), 6(4):32:1– 32:??, October 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3502869.

Paoletti:2018:CLQ

[PPK18] Nicola Paoletti, Andrea Patanè, and Marta Kwiatkowska. Closedloop quantitative verification of rate-adaptive pacemakers. ACM Transactions on Cyber-Physical Systems (TCPS), 2 (4):33:1–33:31, September 2018. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3152767.

Pereyda:2020:CPS

[PRM^{+}20] Christopher Pereyda, Nisha Raghunath, Bryan Minor, Garrett Wilson, Maureen Schmitter-Edgecombe, and Diane J. Cook. Cyber-physical support of daily activities: a robot/smart home partnership. ACM Transactions on Cyber-Physical Systems (TCPS), 4(2): 21:1–21:24, February 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3365225.

Pollicino:2024:PCT

[PSM24] Francesco Pollicino, Dario Stabili, and Mirco Marchetti. Performance comparison of timingbased anomaly detectors for controller area network: a reproducible study. ACM Transactions on Cyber-Physical Systems (TCPS), 8(2):15:1–15:??, April 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3604913.

Panwar:2021:CVP

[PSW^{+}21] Nisha Panwar, Shantanu Sharma, Guoxi Wang, Sharad Mehrotra, and Nalini Venkatasubramanian. CANOPY: a verifiable privacy-preserving token ringbased communication protocol for smart homes. ACM Transactions on Cyber-Physical Systems (TCPS), 5(1):3:1–3:34, January 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3390859.

Pan:2019:MOS

[PXH^{+}19] Chen Pan, Mimi Xie, Song Han, Zhi-Hong Mao, and Jingtong Hu. Modeling and optimization for self-powered non-volatile IoT edge devices with ultralow harvesting power. ACM Transactions on Cyber-Physical Systems (TCPS), 3(3):32:1– 32:26, October 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3324609.

Puthal:2020:UCS

[PYD^{+}20] Deepak Puthal, Laurence T.

Yang, Schahram Dustdar, Zhenyu Wen, Song Jun, Aad van Moorsel, and Rajiv Ranjan. A user-centric security solution for Internet of Things and edge convergence. ACM Transactions on Cyber-Physical Systems (TCPS), 4(3):32:1–32:19, May 2020. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3351882.

Qin:2024:SVU

[QXZ^{+}24] Xin Qin, Yuan Xia, Aditya Zutshi, Chuchu Fan, and Jyotirmoy V. Deshmukh. Statistical verification using surrogate models and conformal inference and a comparison with risk-aware verification. ACM Transactions on Cyber-Physical Systems (TCPS), 8(2):22:1–22:??, April 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3635160.

Rawat:2021:DFA

[RKAG21] Abhimanyu Rawat, Mohammad Khodari, Mikael Asplund, and Andrei Gurtov. Decentralized firmware attestation for invehicle networks. ACM Transactions on Cyber-Physical Systems (TCPS), 5(1):7:1–7:23, January 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3418685.

Roehm:2019:MCC

[ROWA19] Hendrik Roehm, Jens Oehlerking, Matthias Woehrle, and Matthias Althoff. Model conformance for cyber-physical systems: a survey. ACM Transactions on Cyber-Physical Systems (TCPS), 3(3):30:1– 30:26, October 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3306157.

Rosales:2021:CFS

[RP21] Rafael Rosales and Michael Paulitsch. Composable finite state machine-based modeling for quality-of-information-aware cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 5(2):13:1– 13:27, January 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3386244.

Rodionova:2021:LFL

[RPK^{+}21] Alëna Rodionova, Yash Vardhan Pant, Connor Kurtz, Kuk Jang, Houssam Abbas, and Rahul Mangharam. Learning-‘N-Flying: a learning-based, decentralized mission-aware UAS collision avoidance scheme. ACM Transactions on Cyber-Physical Systems (TCPS), 5(4):35:1– 35:26, October 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3447624.

Ramakrishna:2022:EDD

[RRK^{+}22] Shreyas Ramakrishna, Zahra Rahiminasab, Gabor Karsai, Arvind Easwaran, and Abhishek Dubey. Efficient out-of-distribution detection using latent space of β-VAE for cyberphysical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 6(2):15:1–15:??, April 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3491243.

Rathore:2021:TDL

[RSJ21] Heena Rathore, Abhay Samant, and Murtuza Jadliwala. TangleCV: a distributed ledger technique for secure message sharing in connected vehicles. ACM Transactions on Cyber-Physical Systems (TCPS), 5(1): 6:1–6:25, January 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3404500.

Rabadi:2019:RCS

[RTY^{+}19] Dima Rabadi, Rui Tan, David K. Y. Yau, Sreejaya Viswanathan, Hao Zheng, and Peng Cheng. Resilient clock synchronization using power grid voltage. ACM Transactions on Cyber-Physical Systems (TCPS), 3(3):31:1– 31:26, October 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3342048.

Roy:2022:TIA

[RZC^{+}22] Debayan Roy, Licong Zhang, Wanli Chang, Dip Goswami, Birgit Vogel-Heuser, and Samarjit Chakraborty. Tool integration for automated synthesis of distributed embedded controllers. ACM Transactions on Cyber-Physical Systems (TCPS), 6(1): 3:1–3:31, January 2022. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3477499.

Shea-Blymyer:2021:AEF

[SBA21] Colin Shea-Blymyer and Houssam Abbas. Algorithmic ethics: Formalization and verification of autonomous vehicle obligations. ACM Transactions on Cyber-Physical Systems (TCPS), 5(4): 38:1–38:25, October 2021. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3460975.

Sapio:2020:RAW

[SBW20] Adrian Sapio, Shuvra S. Bhattacharyya, and Marilyn Wolf. Runtime adaptation in wireless sensor nodes using structured learning. ACM Transactions on Cyber-Physical Systems (TCPS), 4(4):40:1–40:28, August 2020. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3372153.

Saxena:2020:ISI

[SCB^{+}20] Neetesh Saxena, Alvaro A. Cardenas, Raheem Beyah, Rongxing Lu, Kim-Kwang Raymond Choo, and Yiran Chen. Introduction to the special issue on user-centric security and safety for CPS. ACM Transactions on Cyber-Physical Systems (TCPS), 4(3):29:1–29:2, May 2020. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3392715.

Shrivastava:2021:ISI

[SD21] Aviral Shrivastava and Patricia Derler. Introduction to the special issue on Time for CPS (TCPS). ACM Transactions on Cyber-Physical Systems (TCPS), 5(2):12:1–12:2, January 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3433948.

Sinha:2019:TSR

[SDZV19] Roopak Sinha, Barry Dowdeswell, Gulnara Zhabelova, and Valeriy Vyatkin. TORUS: Scalable requirements traceability for largescale cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 3(2): 15:1–15:25, March 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3203208.

Suntaxi:2021:PSM

[SEB21] Gabriela Suntaxi, Aboubakr Achraf El Ghazi, and Klemens Böhm. Preserving secrecy in mobile social networks. ACM Transactions on Cyber-Physical Systems (TCPS), 5(1):5:1–5:29, January 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3396071.

Sanchez:2022:ECC

[SJT^{+}22] José Manuel Gaspar Sánchez, Nils Jörgensen, Martin Törngren, Rafia Inam, Andrii Berezovskyi, Lei Feng, Elena Fersman, Muhammad Rusyadi Ramli, and Kaige Tan. Edge computing for cyber-physical systems: a systematic mapping study emphasizing trustworthiness. ACM Transactions on Cyber-Physical Systems (TCPS), 6(3):26:1– 26:??, July 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3539662.

Stricker:2023:STC

[SLJ^{+}23] Naomi Stricker, Yingzhao Lian, Yuning Jiang, Colin N. Jones, and Lothar Thiele. Self-triggered control with energy harvesting sensor nodes. ACM Transactions on Cyber-Physical Systems (TCPS), 7(3):20:1–20:??, July 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3597311.

Serru:2023:MCS

[SNBR23] Théo Serru, Nga Nguyen, Michel Batteux, and Antoine Rauzy. Minimal critical sequences in model-based safety and security analyses: Commonalities and differences. ACM Transactions on Cyber-Physical Systems (TCPS), 7(3):17:1–17:??, July 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3593811.

Sheng:2022:PAV

[SPH^{+}22] Shili Sheng, Erfan Pakdamanian, Kyungtae Han, Ziran Wang, John Lenneman, David Parker, and Lu Feng. Planning for automated vehicles with human trust. ACM Transactions on Cyber-Physical Systems (TCPS), 6(4):31:1–31:??, October 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3561059.

Shcherbakov:2022:HDL

[SS22] Maxim Shcherbakov and Cuong Sai. A hybrid deep learning framework for intelligent predictive maintenance of cyberphysical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 6(2):17:1–17:??, April 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3486252.

Sarieddine:2023:ISE

[SST^{+}23] Khaled Sarieddine, Mohammad Ali Sayed, Sadegh Torabi, Ribal Atallah, and Chadi Assi. Investigating the security of EV charging mobile applications as an attack surface. ACM Transactions on Cyber-Physical Systems (TCPS), 7(4):26:1– 26:??, October 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3609508.

Schmidt:2018:CPS

[SSV^{+}18] Mischa Schmidt, Anett Schülke, Alberto Venturi, Roman Kurpatov, and Enrique Blanco Henríquez. Cyber-physical system for energy-efficient stadium operation: Methodology and experimental validation. ACM Transactions on Cyber-Physical Systems (TCPS), 2 (4):25:1–25:26, September 2018. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3140235.

Siegemund:2018:SSP

[ST18] Gerry Siegemund and Volker Turau. A self-stabilizing publish/subscribe middleware for IoT applications. ACM Transactions on Cyber-Physical Systems (TCPS), 2(2):12:1– 12:26, June 2018. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3185509.

Stankovic:2017:RDC

[Sta17] John A. Stankovic. Research directions for cyber physical systems in wireless and mobile healthcare. ACM Transactions on Cyber-Physical Systems (TCPS), 1(1):1:1–1:12, February 2017. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL http:

//dl.acm.org/citation.cfm?
id=2899006.

Samie:2018:DTB

[STB^{+}18] Farzad Samie, Vasileios Tsoutsouras, Lars Bauer, Sotirios Xydis, Dimitrios Soudris, and Jörg Henkel. Distributed tradebased edge device management in multi-gateway IoT. ACM Transactions on Cyber-Physical Systems (TCPS), 2(3): 17:1–17:25, July 2018. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3134842.

Sahani:2023:MLB

[SZCL23] Nitasha Sahani, Ruoxi Zhu, Jin-Hee Cho, and Chen-Ching Liu. Machine learning-based intrusion detection for smart grid computing: a survey. ACM Transactions on Cyber-Physical Systems (TCPS), 7(2):11:1– 11:??, April 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3578366.

Tremel:2019:AFT

[TBKJ19] Edward Tremel, Ken Birman, Robert Kleinberg, and Márk Jelasity. Anonymous, faulttolerant distributed queries for smart devices. ACM Transactions on Cyber-Physical Systems (TCPS), 3(2):16:1– 16:29, March 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3204411.

Trobinger:2022:WCB

[TdAGI^{+}22] Matteo Trobinger, Gabriel de Albuquerque Gleizer, Timofei Istomin, Manuel Mazo, Amy L. Murphy, and Gian Pietro Picco. The wireless control bus: Enabling efficient multi-hop eventtriggered control with concurrent transmissions. ACM Transactions on Cyber-Physical Systems (TCPS), 6(1):4:1–4:29, January 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3485467.

Tamimi:2020:CTI

[THR20] Ali Tamimi, Adam Hahn, and Sandip Roy. Cyber threat impact analysis to air traffic flows through dynamic queue networks. ACM Transactions on Cyber-Physical Systems (TCPS), 4(3):26:1–26:22, March 2020. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/abs/10.1145/3377425.

Tiku:2021:QAD

[TKP21] Saideep Tiku, Prathmesh Kale, and Sudeep Pasricha. Quick-Loc: Adaptive deep-learning for fast indoor localization with mobile devices. ACM Transactions on Cyber-Physical Systems (TCPS), 5(4):44:1–44:30, October 2021. CODEN ????

ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3461342.

Tan:2019:CDN

[TLW^{+}19] Feng Tan, Liansheng Liu, Stefan Winter, Qixin Wang, Neeraj Suri, Lei Bu, Yu Peng, Xue Liu, and Xiyuan Peng. Cross-domain noise impact evaluation for black box two-level control CPS. ACM Transactions on Cyber-Physical Systems (TCPS), 3(1): 2:1–2:25, January 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3226029.

Thakur:2024:CCO

[TMF24] Shailja Thakur, Carlos Moreno, and Sebastian Fischmeister. CANOA: CAN origin authentication through power sidechannel monitoring. ACM Transactions on Cyber-Physical Systems (TCPS), 8(2):13:1– 13:??, April 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571288.

Tan:2017:CLM

[TNY17] Rui Tan, Hoang Hai Nguyen, and David K. Y. Yau. Collaborative load management with safety assurance in smart grids. ACM Transactions on Cyber-Physical Systems (TCPS), 1 (2):12:1–12:27, February 2017. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL http://dl.acm.
                                        org/citation.cfm?id=2823351.

Tu:2021:RSF

[TSC21] Chia-Heng Tu, Qihui Sun, and Hsiao-Hsuan Chang. RAP: a software framework of developing convolutional neural networks for resource-constrained devices using environmental monitoring as a case study. ACM Transactions on Cyber-Physical Systems (TCPS), 5 (4):43:1–43:28, October 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3472612.

Tiausas:2023:HHP

[TYT^{+}23] Francis Tiausas, Keiichi Yasumoto, Jose Paolo Talusan, Hayato Yamana, Hirozumi Yamaguchi, Shameek Bhattacharjee, Abhishek Dubey, and Sajal K. Das. HPRoP: Hierarchical privacy-preserving route planning for smart cities. ACM Transactions on Cyber-Physical Systems (TCPS), 7(4):27:1– 27:??, October 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3616874.

Valente:2019:ISV

[VBV^{+}19] Junia Valente, Kanchan Bahirat, Kelly Venechanos, Alvaro A. Cardenas, and Prabhakaran Balakrishnan. Improving the security of visual challenges. ACM Transactions on Cyber-Physical Systems (TCPS), 3(3): 34:1–34:26, October 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3331183.

Vinitsky:2023:OMA

[VLPB23] Eugene Vinitsky, Nathan Lichtlé, Kanaad Parvate, and Alexandre Bayen. Optimizing mixed autonomy traffic flow with decentralized autonomous vehicles and multi-agent reinforcement learning. ACM Transactions on Cyber-Physical Systems (TCPS), 7(2):13:1–13:??, April 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3582576.

VanDeHoef:2020:PFD

[VMDJ20] Sebastian Van De Hoef, Jonas Mårtensson, Dimos V. Dimarogonas, and Karl Henrik Johansson. A predictive framework for dynamic heavy-duty vehicle platoon coordination. ACM Transactions on Cyber-Physical Systems (TCPS), 4(1): 5:1–5:25, January 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3299110.

Varadharajan:2024:TES

[VTK24] Vijay Varadharajan, Uday Tupakula, and Kallol Krishna Karmakar. Techniques for enhancing security in industrial control systems. ACM Transactions on Cyber-Physical Systems (TCPS), 8(1):6:1–6:??, January 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3630103.

Waga:2022:MBM

[WAH22] Masaki Waga, Étienne André, and Ichiro Hasuo. Modelbounded monitoring of hybrid systems. ACM Transactions on Cyber-Physical Systems (TCPS), 6(4):30:1–30:??, October 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3529095.

Wang:2024:GDC

[WCZ^{+}24] Ruihang Wang, Zhiwei Cao, Xin Zhou, Yonggang Wen, and Rui Tan. Green data center cooling control via physics-guided safe reinforcement learning. ACM Transactions on Cyber-Physical Systems (TCPS), 8(2):19:1– 19:??, April 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3582577.

Wen:2021:CLQ

[WD21] Jianjun Wen and Waltenegus Dargie. Characterization of link quality fluctuation in mobile wireless sensor networks. ACM Transactions on Cyber-Physical Systems (TCPS), 5(3): 28:1–28:24, July 2021. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3448737.

Wrana:2022:OTS

[WEL^{+}22] Michael Maximilian Wrana, Marwa Elsayed, Karim Lounis, Ziad Mansour, Steven Ding, and Mohammad Zulkernine. OD1NF1ST: True skip intrusion detection and avionics network cyber-attack simulation. ACM Transactions on Cyber-Physical Systems (TCPS), 6 (4):33:1–33:??, October 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3551893.

Wang:2017:ESH

[WGH^{+}17] Hongwei Wang, Yunlong Gao, Shaohan Hu, Shiguang Wang, Renato Mancuso, Minje Kim, Poliang Wu, Lu Su, Lui Sha, and Tarek Abdelzaher. On exploiting structured human interactions to enhance sensing accuracy in cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 1(3): 16:1–16:19, July 2017. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3064006.

Wang:2019:RTM

[WGL19] Chao Wang, Christopher Gill, and Chenyang Lu. Realtime middleware for cyberphysical event processing. ACM Transactions on Cyber-Physical Systems (TCPS), 3(3):29:1– 29:25, October 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3218816.

Wang:2018:RRC

[WHS18] Xiaofeng Wang, Naira Hovakimyan, and Lui Sha. RSimplex: a robust control architecture for cyber and physical failures. ACM Transactions on Cyber-Physical Systems (TCPS), 2(4):27:1–27:26, September 2018. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3121428.

Weber:2020:GFR

[WJL^{+}20] Matthew Weber, Baihong Jin, Gil Lederman, Yasser Shoukry, Edward A. Lee, Sanjit Seshia, and Alberto Sangiovanni-Vincentelli. Gordian: Formal reasoning-based outlier detection for secure localization. ACM Transactions on Cyber-Physical Systems (TCPS), 4(4): 43:1–43:27, August 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3386568.

Wu:2018:DSA

[WKS18] Qiang Wu, T. John Koo, and Yoshihiko Susuki. Dynamic security analysis of power systems by a sampling-based algorithm. ACM Transactions on Cyber-Physical Systems (TCPS), 2(2):

10:1–10:26, June 2018. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3208093.

Wan:2019:PLK

[WLA19] Jiang Wan, Anthony Lopez, and Mohammad Abdullah Al Faruque. Physical layer key generation: Securing wireless communication in automotive cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 3(2): 13:1–13:26, March 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3140257.

Wang:2022:RTC

[WLTW22] Rongrong Wang, Duc Van Le, Rui Tan, and Yew-Wah Wong. Real-time cooling power attribution for co-located data center rooms with distinct temperatures and humidities. ACM Transactions on Cyber-Physical Systems (TCPS), 6(1):7:1–7:28, January 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3494578.

Wang:2018:LTE

[WMC18] Ping Wang, Meng Ma, and Chao-Hsien Chu. Longterm event processing over data streams in cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 2(2):14:1–14:23, June 2018. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/abs/10.1145/3204412.

Wang:2024:TSA

[WSH^{+}24] Jiangwei Wang, Lili Su, Songyang Han, Dongjin Song, and Fei Miao. Towards safe autonomy in hybrid traffic: Detecting unpredictable abnormal behaviors of human drivers via information sharing. ACM Transactions on Cyber-Physical Systems (TCPS), 8(2):16:1–16:??, April 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3616398.

Wu:2019:CYC

[WTZ^{+}19] Tingmin Wu, Lihong Tang, Rongjunchen Zhang, Sheng Wen, Cecile Paris, Surya Nepal, Marthie Grobler, and Yang Xiang. Catering to your concerns: Automatic generation of personalised security-centric descriptions for Android apps. ACM Transactions on Cyber-Physical Systems (TCPS), 3(4): 36:1–36:21, October 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3317699.

Wang:2022:RAI

[WWC22] Michael I.-C. Wang, Charles H.- P. Wen, and H. Jonathan Chao. Roadrunner+: an autonomous intersection management cooperating with connected autonomous vehicles and pedestrians with spillback considered. ACM Transactions on Cyber-Physical Systems (TCPS), 6(1): 5:1–5:29, January 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3488246.

Wang:2019:ILD

[WWL^{+}19] Tian Wang, Wenhua Wang, Anfeng Liu, Shaobin Cai, and Jiannong Cao. Improve the localization dependability for cyber-physical applications. ACM Transactions on Cyber-Physical Systems (TCPS), 3(1): 6:1–6:21, January 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3140240.

Wang:2020:EMF

[WWW20] Ding Wang, Ping Wang, and Chenyu Wang. Efficient multifactor user authentication protocol with forward secrecy for real-time data access in WSNs. ACM Transactions on Cyber-Physical Systems (TCPS), 4(3): 30:1–30:26, March 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3325130.

Wang:2019:DTT

[WYW^{+}19] Xiaokang Wang, Laurence T.

Yang, Yihao Wang, Xingang Liu, Qingxia Zhang, and M. Jamal Deen. A distributed tensortrain decomposition method for cyber-physical-social services. ACM Transactions on Cyber-Physical Systems (TCPS), 3(4): 35:1–35:15, October 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3323926.

Wu:2017:DRW

[WZ17] Jianjia Wu and Wei Zhao. Design and realization of WInternet: From net of things to Internet of Things. ACM Transactions on Cyber-Physical Systems (TCPS), 1(1):2:1–2:12, February 2017. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL http:

                                        //dl.acm.org/citation.cfm?
                                        id=2872332.

Wei:2019:ISI

[WZR^{+}19] Tongquan Wei, Junlong Zhou, Rajiv Ranjan, Isaac Triguero, Huafeng Yu, Chun Jason Xue, and Schahram Dustdar. Introduction to the special issue on human-interaction-aware data analytics for cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 3(4):35e:1–35e:2, October 2019. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/3344260.

Wang:2019:DTS

[WZS^{+}19] Chang Wang, Yongxin Zhu, Weiwei Shi, Victor Chang, P. Vijayakumar, Bin Liu, Yishu Mao, Jiabao Wang, and Yiping Fan. A dependable time series analytic framework for cyber-physical systems of IoT-based smart grid. ACM Transactions on Cyber-Physical Systems (TCPS), 3(1): 7:1–7:18, January 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3145623.

Wu:2020:ADB

[WZZ^{+}20] Di Wu, Hanlin Zhu, Yongxin Zhu, Victor Chang, Cong He, Ching-Hsien Hsu, Hui Wang, Songlin Feng, Li Tian, and Zunkai Huang. Anomaly detection based on RBM-LSTM neural network for CPS in advanced driver assistance system. ACM Transactions on Cyber-Physical Systems (TCPS), 4(3): 27:1–27:17, May 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3377408.

Xie:2019:HIA

[XBW^{+}19] Guoqi Xie, Yang Bai, Wei Wu, Yanwen Li, Renfa Li, and Keqin Li. Human-interaction-aware adaptive functional safety processing for multi-functional automotive cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 3(4):39:1–39:25, October 2019. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3337931.

Xiao:2019:ILD

[XCL^{+}19] Fu Xiao, Jing Chen, Zhetao Li, Haiping Huang, and Lijuan Sun. Improved LDA dimension reduction based behavior learning with commodity WiFi for cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 3(4): 38:1–38:19, October 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3342219.

Xie:2019:RCA

[XZA^{+}19] Guoqi Xie, Gang Zeng, Jiyao An, Renfa Li, and Keqin Li. Resource-cost-aware faulttolerant design methodology for end-to-end functional safety computation on automotive cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 3(1):4:1– 4:27, January 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3162052.

Yip:2018:TEC

[YAR^{+}18] Eugene Yip, Sidharta Andalam, Partha S. Roop, Avinash Malik, Mark L. Trew, Weiwei Ai, and Nitish Patel. Towards the emulation of the cardiac conduction system for pacemaker validation. ACM Transactions on Cyber-Physical Systems (TCPS), 2 (4):32:1–32:26, September 2018. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3134845.

Yang:2022:FIV

[YBKJ22] Xiaodong Yang, Omar Ali Beg, Matthew Kenigsberg, and Taylor T. Johnson. A framework for identification and validation of affine hybrid automata from input-output traces. ACM Transactions on Cyber-Physical Systems (TCPS), 6(2):13:1– 13:??, April 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

https://dl.acm.org/doi/10.
1145/3470455.

Yang:2019:CBC

[YFS^{+}19] Chaoqun Yang, Li Feng, Zhiguo Shi, Rongxing Lu, and Kim-Kwang Raymond Choo. A crowdsensing-based cyber-physical system for drone surveillance using random finite set theory. ACM Transactions on Cyber-Physical Systems (TCPS), 3(4): 42:1–42:22, October 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3342049.

Yan:2022:COA

[YHL^{+}22] Junjie Yan, Kevin Huang, Kyle Lindgren, Tamara Bonaci, and Howard J. Chizeck. Continuous operator authentication for teleoperated systems using hidden Markov models. ACM Transactions on Cyber-Physical Systems (TCPS), 6(1):6:1–6:25, January 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3488901.

Yang:2024:MBD

[YKDL24] Yahan Yang, Ramneet Kaur, Souradeep Dutta, and Insup Lee. Memory-based distribution shift detection for learning enabled cyber-physical systems with statistical guarantees. ACM Transactions on Cyber-Physical Systems (TCPS), 8(2): 21:1–21:??, April 2024. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3643892.

Yu:2023:DDP

[YL23] Handi Yu and Xin Li. Datadriven parameterized corner synthesis for efficient validation of perception systems for autonomous driving. ACM Transactions on Cyber-Physical Systems (TCPS), 7(2):9:1–9:??, April 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3571286.

Yuan:2022:DDC

[YMH^{+}22] Yukun Yuan, Meiyi Ma, Songyang Han, Desheng Zhang, Fei Miao, John A. Stankovic, and Shan Lin. DeResolver: a decentralized conflict resolution framework with autonomous negotiation for smart city services. ACM Transactions on Cyber-Physical Systems (TCPS), 6 (4):29:1–29:??, October 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3529096.

Yu:2022:EER

[YNL^{+}22] Ping Yu, Wei Ni, Ren Ping Liu, Zhaoxin Zhang, Hua Zhang, and Qiaoyan Wen. Efficient encrypted range query on cloud platforms. ACM Transactions on Cyber-Physical Systems (TCPS), 6(3):27:1–27:??, July 2022. CODEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3548657.

Yan:2020:TOV

[YS20] Li Yan and Haiying Shen.

TOP: Optimizing vehicle driving speed with vehicle trajectories for travel time minimization and road congestion avoidance. ACM Transactions on Cyber-Physical Systems (TCPS), 4(2): 17:1–17:25, February 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3362162.

Yan:2021:UGT

[YS21] Li Yan and Haiying Shen. Utilizing game theory to optimize inmotion wireless charging service efficiency for electric vehicles. ACM Transactions on Cyber-Physical Systems (TCPS), 5(2): 19:1–19:26, January 2021. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3430194.

Yoon:2019:SAP

[YWM^{+}19] Hyung-Jin Yoon, Christopher Widdowson, Thiago Marinho, Ranxiao Frances Wang, and Naira Hovakimyan. Socially aware path planning for a flying robot in close proximity of humans. ACM Transactions on Cyber-Physical Systems (TCPS), 3(4):41:1– 41:24, October 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3341570.

Yu:2019:PPD

[YWZ^{+}19] Jiahui Yu, Kun Wang, Deze Zeng, Chunsheng Zhu, and Song Guo. Privacy-preserving data aggregation computing in cyber-physical social systems. ACM Transactions on Cyber-Physical Systems (TCPS), 3(1): 8:1–8:23, January 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3145625.

Yong:2018:SDI

[YZF18] Sze Zheng Yong, Minghui Zhu, and Emilio Frazzoli. Switching and data injection attacks on stochastic cyber-physical systems: Modeling, resilient estimation, and attack mitigation. ACM Transactions on Cyber-Physical Systems (TCPS), 2 (2):9:1–9:2, June 2018. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3204439.

Zhao:2018:PSIa

[ZA18a] Wei Zhao and Tarek Abdelzaher. Preface to the special issue: Toward an efficient and effective Internet of Things for cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 2 (2):7:1–7:2, June 2018. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3203210.

Zhao:2018:PSIb

[ZA18b] Wei Zhao and Tarek Abdelzaher. Preface to the special issue: Toward an efficient and effective Internet of Things for cyber-physical systems (Part II). ACM Transactions on Cyber-Physical Systems (TCPS), 2 (3):15:1–15:2, July 2018. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3203213.

Zhang:2020:HMT

[ZDY^{+}20] Zhenyong Zhang, Ruilong Deng, David K. Y. Yau, Peng Cheng, and Jiming Chen. On hiddenness of moving target defense against false data injection attacks on power grid. ACM Transactions on Cyber-Physical Systems (TCPS), 4(3): 25:1–25:29, March 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3372751.

Zoppi:2021:MLI

[ZGAB21] Tommaso Zoppi, Mohamad Gharib, Muhammad Atif, and Andrea Bondavalli. Metalearning to improve unsupervised intrusion detection in cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 5(4):42:1– 42:27, October 2021. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3467470.

Zhang:2019:NST

[ZHZ19] Desheng Zhang, Tian He, and Fan Zhang. Nationalscale traffic model calibration in real time with multisource incomplete data. ACM Transactions on Cyber-Physical Systems (TCPS), 3(2):21:1– 21:26, March 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3300186.

Zhong:2021:EST

[ZJP21] Bingzhuo Zhong, Claudius Jordan, and Julien Provost. Extending signal temporal logic with quantitative semantics by intervals for robust monitoring of cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 5(2): 16:1–16:25, January 2021. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3377868.

Zheng:2020:DAD

[ZLSZ20] Bowen Zheng, Chung-Wei Lin, Shinichi Shiraishi, and Qi Zhu. Design and analysis of delaytolerant intelligent intersection management. ACM Transactions on Cyber-Physical Systems (TCPS), 4(1):3:1– 3:27, January 2020. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3300184.

Zoppi:2023:SMR

[ZMC^{+}23] Tommaso Zoppi, Innocenzo Mungiello, Andrea Ceccarelli, Alberto Cirillo, Lorenzo Sarti, Lorenzo Esposito, Giuseppe Scaglione, Sergio Repetto, and Andrea Bondavalli. Safe maintenance of railways using COTS mobile devices: The remote worker dashboard. ACM Transactions on Cyber-Physical Systems (TCPS), 7(4):25:1–25:??, October 2023. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3607193.

Zimmerling:2017:ART

[ZMK^{+}17] Marco Zimmerling, Luca Mottola, Pratyush Kumar, Federico Ferrari, and Lothar Thiele. Adaptive real-time communication for wireless cyberphysical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 1(2):8:1–8:29, February 2017. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL http:

                                        //dl.acm.org/citation.cfm?
                                        id=3012005.

Zhang:2019:DDC

[ZYCL19] Qingchen Zhang, Laurence T. Yang, Zhikui Chen, and Peng Li. Dependable deep computation model for feature learning on big data in cyber-physical systems. ACM Transactions on Cyber-Physical Systems (TCPS), 3(1): 11:1–11:17, January 2019. CO-DEN ???? ISSN 2378-962X (print), 2378-9638 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3110218.

Zhao:2024:COA

[ZZC^{+}24] Kejing Zhao, Zhiyong Zhang, Kim-Kwang Raymond Choo, Zhongya Zhang, and Tiantian Zhang. A combinatorial optimization analysis method for detecting malicious industrial Internet attack behaviors. ACM Transactions on Cyber-Physical Systems (TCPS), 8(1):8:1–8:??, January 2024. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3637554.

Zhang:2017:HMI

[ZZZ^{+}17a] Desheng Zhang, Juanjuan Zhao, Fan Zhang, Tian He, Haengju Lee, and Sang H. Son. Heterogeneous model integration for multi-source urban infrastructure data. ACM Transactions on Cyber-Physical Systems (TCPS), 1(1):4:1–4:26, February 2017. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL http:

//dl.acm.org/citation.cfm?
id=2967503.

Zhang:2017:LMT

[ZZZ^{+}17b] Desheng Zhang, Juanjuan Zhao, Fan Zhang, Ruobing Jiang, Tian He, and Nikos Papanikolopoulos. Last-mile transit service with urban infrastructure data. ACM Transactions on Cyber-Physical Systems (TCPS), 1(2):6:1–6:26, February 2017. CODEN ???? ISSN 2378-962X (print), 2378-9638 (electronic). URL http:

//dl.acm.org/citation.cfm?
id=2823326.